### v1

# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

Abstract
Process mapping asks to assign vertices of a task graph to processing elements of a supercomputer such that the computational workload is balanced while the communication cost is minimized. Motivated by the recent success of GPU-based graph partitioners, we propose two GPU-accelerated algorithms for this optimization problem.
The first algorithm employs hierarchical multisection, which partitions the task graph alongside the hierarchy of the supercomputer. The method utilizes GPU-based graph partitioners to accelerate the mapping process. The second algorithm integrates process mapping directly into the modern multilevel graph partitioning pipeline. Vital phases like coarsening and refinement are accelerated by exploiting the parallelism of GPUs.
In our experiments, both methods achieve speedups exceeding 300 when compared to state-of-the-art CPU-based algorithms. The first algorithm has, on average, about 10 percent greater communication costs and thus remains competitive to CPU algorithms. The second approach is much faster, with a geometric mean speedup of 77.6 and peak speedup of 598 at the cost of lower solution quality. To our knowledge, these are the first GPU-based algorithms for process mapping.
I Introduction
Parallel high-performance applications rely on the efficient execution of large numbers of tasks across massively parallel architectures. This involves placing individual tasks on the Processing Elements (PEs) of the underlying hardware architecture such that the computational workload is evenly balanced. However, load balancing alone is not sufficient. In parallel applications, the individual tasks usually communicate with each other, i.e., they send and receive data to and from each other. The cost of this communication heavily depends on the location of the tasks in hardware topology. Modern supercomputers are typically organized hierarchically, for example into processors, nodes, racks and islands. The communication cost between two PEs, therefore, depends on their relative locations within this hierarchy. Two PEs on the same processor communicate data much more quickly than two PEs which are located on different islands. Thus, in addition to balancing the workload, it is advantageous to minimize the total communication cost by placing highly interactive tasks close to each other. This optimization problem is known as the general process mapping problem (GPMP).
Optimally solving the GPMP is $\mathrm{NP}$-hard [39] as it involves solving the well-known quadratic assignment problem (QAP) [9]. The QAP asks for a set of $n$ facilities and $n$ locations to minimize the total cost of assigning each facility to one location, where the cost depends on both the flow between facilities and the distance between their locations. Since the QAP is strongly $\mathrm{NP}$-hard, i.e., no constant factor approximation exists unless $\mathrm{P}=\mathrm{NP}$, its hardness transfers to the GPMP. The hardness is also reflected in practice as no meaningful instance with $n>20$ can be solved in a reasonable time [9]. Moreover, the same hardness result arises from graph partitioning (GP), where the goal is to partition a graph while minimizing the number of edges between them. Graph partitioning is $\mathrm{NP}$-complete and also admits no constant-factor approximation unless $\mathrm{P}=\mathrm{NP}$.
The GPMP is commonly solved using either the two-phase approach or via integrated mapping. In the two-phase approach, the task set is first partitioned into as many (almost) equally sized sets as there are PEs in the supercomputer. While partitioning the task set, the necessary communication between the resulting sets is minimized. In the second phase, the subsets are mapped to the PEs such that the total cost of communication is minimized, taking the distances between the PEs into account. The first phase is an instance of GP, while second phase is a QAP instance. The two-phase approach is often easier to realize as it can fall back to two well-studied problems, i.e., graph partitioning the QAP. The algorithm SharedMap [44] currently offers state-of-the-art solution quality by utilizing hierarchical multisection, a special variant of the two-phase approach.
The integrated mapping approach combines both phases into a single process by incorporating the mapping objective directly into the graph partitioning pipeline. Rather than minimizing the number of edges crossing between the subsets, as in classical graph partitioning, it directly minimizes the total cost of communication. Typically, this is achieved using the modern multilevel framework, which involves coarsening the input graph to a smaller representation, computing an initial mapping, and then refining it during uncoarsening. The approach is more complex to realize as the whole graph partitioning pipeline has to be adopted to the new objective function. However, this also enables function specific optimization, which result in an overall faster execution time. As a result, integrated mapping often achieves faster runtimes compared to the two-phase approach.
Recently, Graphic Processing Units (GPUs) have been utilized to tackle graph partitioning. GPUs provide massive parallelism that can significantly reduce runtime, when exploited effectively. However, leveraging thousands of GPU cores is challenging due to the irregularity of graphs and lack of straightforward parallelization strategies. Despite this, recent GPU-based algorithms like Jet [18] and G-kway [31] have demonstrated promising results. They deliver solution qualities comparable to the serial CPU-based partitioner Metis [27], while being up to 20 times faster. Since both the two-phase and integrated mapping approach rely on graph partitioning, these results should transfer to GPMP.
[…]
II Preliminaries
II-A Concepts and Notations
Let $G=(V,E)$ with $n=|V|$ and $m=|E|$ be an undirected graph, and let $c:V\mapsto\mathbb{R}$ and $\omega:E\mapsto\mathbb{R}$ be vertex and edge weights. Their extension to sets is $c(V^{\prime})=\sum_{v\in V^{\prime}}c(v)$ and $\omega(E^{\prime})=\sum_{e\in E^{\prime}}\omega(e)$ for vertex sets $V^{\prime}$ and edge sets $E^{\prime}$.
The graph partitioning problem (GPP) asks for a partition of the vertex set $V$ into $k$ equally sized sets. More formally let $k\in\mathbb{N}$ be the desired number of blocks and let $\epsilon\geq 0$ be a hyper-parameter called imbalance. GPP then asks for a partition $V=V_{1}\cup\ldots\cup V_{k}$ such that $\forall i\neq j:V_{i}\cap V_{j}=\emptyset$ and each $V_{i}$ adheres to the balance constraint, that is $c(V_{i})\leq L_{\max}\coloneqq(1+\epsilon)\frac{c(V)}{k}$. Such a partition is called a $k$-way partition of $G$. Let $E_{ij}=\{\{u,v\}\in E\mid u\in V_{i},v\in V_{j}\}$ be the edges crossing between block $V_{i}$ and $V_{j}$. The edge-cut of a partition is the total weight of edges crossing between blocks, i.e., $\sum_{i<j}\omega(E_{ij})$. The goal of GP is to determine a balanced $k$-way partition that minimizes the edge-cut.
A parallel application consists of $n$ weighted tasks which communicate with each other. This is captured by the communication matrix $\mathcal{C}\in\mathbb{R}^{n\times n}$, where an entry $\mathcal{C}_{ij}$ denotes the communication volume between tasks $i$ and $j$. An entry $\mathcal{C}_{ij}=0$ denotes that the tasks do not communicate with each other. Most modern parallel applications only have a very sparse communication pattern between the tasks, which is why we model the communication matrix $\mathcal{C}$ as a task graph $G_{\mathcal{C}}$. For each non-zero entry $\mathcal{C}_{ij}$ the task graph has one forward and backward edge with weight $\mathcal{C}_{ij}$.
Let $k$ be the number of PEs on a supercomputer. The cost of communication between the PEs is captured by the hardware topology matrix $\mathcal{D}\in\mathbb{R}^{k\times k}$. An entry $\mathcal{D}_{xy}$ denotes a multiplicative cost factor that is applied to communication between PEs $x$ and $y$. The cost of task $i$ and $j$ communicating, if placed on PEs $x$ and $y$ respectively, is $\mathcal{C}_{ij}\mathcal{D}_{xy}$.
The focus of this work is to tackle the general process mapping problem. It asks to assign the vertices of the task graph $G_{\mathcal{C}}$ to PEs of the supercomputer such that the amount of work on each PE is balanced and the total cost of communication is minimized. More formally, we ask for a mapping $\Pi:[n]\mapsto[k]$ such that $J(C,D,\Pi)\coloneqq\sum_{i,j}\mathcal{C}_{ij}\mathcal{D}_{\Pi(i)\Pi(j)}$ is minimized, while adhering to the balance constraint.
In the hierarchical process mapping problem (HPMP), a special variant of the GPMP, the hardware topology of the supercomputer is described by a homogeneous hierarchy $H=a_{1}:\ldots:a_{\ell}$. It denotes that each processor has $a_{1}$ PEs, each node has $a_{2}$ processors, each rack has $a_{3}$ nodes, and so forth. In total, the system has $k=\prod_{1\leq\ell}a_{i}$ PEs. Corresponding to the hierarchy, a distance $D=d_{1}:\ldots:d_{\ell}$ is given which denotes the cost factor between the PEs. Two PEs on the same processor have a cost factor of $d_{1}$, two PEs on the same node but on different processors have a cost factor of $d_{2}$ and so forth.
In GPMP we typically assume $n>k$, i.e., we have (considerably) more tasks than PEs. If $n=k$ the problem is known as the one-to-one process mapping problem (OPMP) which is identical to the quadratic assignment problem (QAP). The QAP is commonly solved using refactorization and linearization techniques [3, 2, 21], of its corresponding mixed integer linear programming formulation [30]. We refer the reader to [9, 4, 32] for comprehensive surveys. The problems GPP, GPMP, HPMP and QAP are all NP-hard problems [17, 39] and no constant factor approximation guarantee exists unless P = NP.
II-B Multilevel Graph Partitioning
Modern graph partitioning relies on the multilevel approach, which works in three phases, coarsening, partitioning and uncoarsening with refinement. In the coarsening phase a succession of smaller graphs $G,G_{1},\ldots,G_{t}$ is computed, with the goal that a smaller graph $G_{i+1}$ retains the same overall structure as $G_{i}$. A “good” partition of $G_{i+1}$ would then also resemble a “good” partition for $G_{i}$. Coarsening is applied until the coarsest graph has less than $ck$ vertices, where $c$ is a tunable constant. To coarsen a graph, one usually applies a matching or a clustering algorithm. The matching algorithm rates each edge and then determines a maximum weighted matching. Edge contraction is then applied on the determined matching. Edge contraction on an edge $\{u,v\}$ is applied by replacing vertices $u$ and $v$ with a new vertex $w$ with weight $c(w)=c(u)+c(v)$. Former neighbors of $u$ and $v$ are connected to $w$ with the same edge weight. Parallel edges are fused and their weight is summed. A clustering algorithm first determines clusters in the graph, and then contracts each cluster into one super-vertex. The vertex weights and parallel edges are handled in the same way as in the matching.
In the initial partitioning phase, the coarsest graph $G_{t}$ is partitioned. Since $G_{t}$ is very small, expensive but high-quality algorithms, such as recursive bisection or direct $k$-way partitioning, can be applied to minimize edge-cut.
The uncoarsening and refinement phase iteratively uncontracts the graphs $G_{i+1}$ into $G_{i}$ by undoing the contractions. After each uncontraction, local search algorithms are applied to improve the objective. The algorithms try to move vertices across the blocks to minimize the edge-cut, while respecting the balance constraint. Widely used refinement algorithms are the Fiduccia-Mattheyses (FM) algorithm [16], label propagation [38] and flow-based refinement [40]. Notable serial graph partitioners are Metis [27], KaFFPa [40], Jostle [48], Scotch [36] and KaHyPar [41].
III Related Work
Process mapping is closely related to graph partitioning, which has already seen a tremendous amount of research over the last few decades. We already gave a brief overview in Section II-B and refer the reader to [6, 10, 8, 42] for more comprehensive surveys. We will here give an overview of Jet [18], a GPU-based graph partitioning (Section III-A), and of process mapping (Section III-B).
III-A GPU Multilevel Graph Partitioning
Gilbert et al. [18] present Jet, a GPU-accelerated multilevel graph partitioner designed to achieve high partition quality and fast runtimes. Their main contribution includes a novel GPU-parallel refinement algorithm based on label propagation [38]. They implemented their algorithm in the performance-portable Kokkos framework [14], enabling execution on GPUs and multi-core CPUs with a single code base.
Matching. Jet utilizes a two-hop matching [29], originally developed for Mt-Metis [28], a shared-memory graph partitioner. The algorithm first applies a standard heavy-edge matching and only if more than 25% of vertices remain unmatched a two hop-matching is applied. A heavy edge-matching determines for each vertex its heaviest edge and two neighbors are matched if they agree on their heaviest edge. The two-hop matching can be split into three categories: leaf, twin and relative vertices. Leaf vertices have degree one, two twin vertices share the same neighborhood and relative vertices share at least one common neighbor. To identify twin vertices the neighborhoods of all vertices are hashed and sorted, so vertices with the same neighborhood can easily be found. To accelerate the matching of relatives, special matchmaker vertices with small degree are used. The two-hop matching first matches leaf vertices that share a neighbor, then twin vertices and in the end relative vertices, if necessary.
Coarsening. To coarsen the graph, a fine-grained per vertex hashing algorithm is deployed. First the maximum neighborhood size of each vertex in the coarse graph is overestimated. Then a prefix-sum over the sizes is determined, assigning each vertex a subarray to store its neighbors. Each vertex hashes its new neighboring vertices and inserts them and their weight into their respective subarray. In the end, the edges of all vertices are extracted into the classic Compressed Sparse Row (CSR) format.
Initial Partitioning. After coarsening to 4k4k8k8k vertices, the partitioning of the coarsest graph is delegated to Metis [27] on the CPU. The transfer overhead between CPU and GPU is negligible due to the small size of the coarsest graph.
Refinement. The novel improvement of Jet is its refinement algorithm, which alternates between a highly parallelized label propagation [38] algorithm and two rebalancing strategies. A standard serial label propagation algorithm visits the vertices in random order. A vertex is moved to a neighboring block if it improves the edge-cut and adheres to the balance constraint. This is often repeated in multiple rounds to further refine the partition. However, the local search can get stuck in local optima as only positive moves are allowed. Additional problems arise with parallelization. Moving multiple vertices at once makes it difficult to maintain the balance constraint. Additionally, while each individual vertex move might have a positive impact, moving adjacent vertices in parallel can worsen the solution quality.
Gilbert et. al [18] tackle these problems by splitting the refinement into two parts, first unconstrained label propagation and afterwards rebalancing. The unconstrained label propagation starts by determining for each vertex $v$ the gain $F(v)$ (decrease in edge-cut) of moving it in an adjacent block. All potential vertex moves are then passed through the first filter $F(v)\geq 0\lor-F(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. Here, $\operatorname{conn}(v,b)$ is the sum of edge weights from vertex $v$ to block $b$. To pass the first filter a move needs to be non-negative (fist condition) or its increase in edge-cut must be comparatively small (second condition). The constant $c\in[0,1]$ controls the number of negative gain moves. These moves are allowed to escape local minima. All moves that pass the first filter are collected in a list $\mathcal{X}$, which is implicitly sorted decreasingly by gain. For each vertex $v\in\mathcal{X}$, the algorithm then recomputes the gain $F_{2}(v)$ under an approximate future partition state: if a neighbor $u$ appears earlier in the ordering, it is assumed to have already moved to its new block. In a second filter only the vertices $v\in\mathcal{X}$ with a non-negative gain $F_{2}(v)$ are moved to their respective destination. To prevent oscillation each vertex that was moved is locked in the next iteration of unconstrained label propagation.
Rebalancing. Since the unconstrained label propagation phase ignores the balance constraint, an additional rebalancing phase is required. Jet implements two similar strategies, weak and strong rebalancing. Weak rebalancing determines for each vertex in an oversized partition $a$ the loss of moving it to a valid destination block. Valid destination blocks are blocks $b$ with $c(b)\leq\sigma\leq L_{\max}$. The upper bound $\sigma$ is used instead of $L_{\max}$ to create a small buffer zone, so the block is not overloaded after only one move. For each oversized block $a$ its vertices are approximately sorted increasingly by loss in a list $L_{a}$. Then a prefix $L_{a}^{*}\subseteq L_{a}$ is determined such that moving all vertices in $L_{a}^{*}$ will balance block $a$. This minimizes the loss to balance a specific block, however, the destination blocks might become oversized, so another iteration might be necessary. Strong rebalancing reverses the process. Instead of oversized blocks pushing their vertices out, underloaded blocks will pull vertices from oversized blocks. They will only pull so many vertices such they themselves will not become oversized. This often achieves a balanced partition in just one iteration, however the loss is typically greater as in weak rebalancing. In practice, Jet applies two iterations of weak rebalancing followed by one iteration of strong rebalancing if necessary.
In total, Jet combines its unconstrained label propagation and rebalancing in multiple iterations and keeps the best solution encountered. The algorithm offers comparable solution quality to the well established CPU-based graph partitioner Metis [27], while being up to 20 times faster. Other GPU graph partitioners are Goodarzi et al. [19], Sphynx [1] and G-kway [31].
III-B Process Mapping
Similar to graph partitioning, process mapping has also seen a tremendous amount of research, and we refer the reader to Hoefler et. al [25] and Pellegrini [37] for broader overviews. Among the first to implement process mapping, was Hatazaki [22] who used graph partitioning to map MPI processes onto hardware topologies. Träff [46] and Yu et. al  [49] implemented mapping algorithms for special hardware, namely for the NEC SX-series of parallel vector computers and the Blue Gene/L Supercomputer, respectively.
From a theoretical perspective it seems straightforward that minimizing $J(C,D,\Pi)$ will lead to a more performant algorithm. However, the questions remains if this is also true in practice. Brandfass et. al [7] showed that optimizing the communication cost indeed leads to performance improvements in the context of Computational Fluid Dynamics. Manwade et. al [33] show in a more general context that process mapping algorithms improve the performance of MPI applications.
There are two main methods to solve process mapping, the two-phase approach and integrated mapping. The two-phase approach first leverages standard graph partitioning to determine a balanced $k$-way partition with small edge-cut of the task graph $G_{\mathcal{C}}$. The second phase then determines a 1-to-1 mapping of blocks to the PEs that minimizes the communication cost. Müller-Merbach [35] presented a greedy algorithm to obtain an initial mapping, that served as a basis for many future algorithms. After the mapping phase a potential third refinement phase can be used to determine individual vertex moves that decrease the communication cost. Heider [23] introduced a simple swap algorithm that considers all possible $\mathcal{O}(n^{2})$ swaps. This method was iteratively improved by Brandfass et. al [7] by avoiding redundant swaps, and Schulz and Träff [43], by utilizing better data structures and restricting the search space. Notable algorithms using the two-phase approach are GlobalMultisection [47] and SharedMap [44].
GlobalMultisection [47] employs a method called hierarchical multisection inspired by KaFFPa-Map [43]. KaFFPa-Map would first create a balanced $k$-way partition and then create a communication model graph $G_{\mathbb{M}}$. This graph has $k$ vertices and edges connecting them if the corresponding $k$-way partition has edges between the blocks. $G_{\mathcal{M}}$ is then perfectly partitioned alongside the hierarchy $H=a_{1}:\ldots:a_{\ell}$, first into an $a_{\ell}$-way partition, then each of the blocks into an $a_{\ell-1}$-way partition and so forth. This is recursively repeated until $k$ blocks are obtained which represent the $k$ PEs of the system. The mapping of blocks to PEs, and therefore vertices to PEs, can be determined by following the recursive calls. GlobalMultisection improves the approach by not creating the communication model graph $G_{\mathbb{M}}$, but instead directly determining an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ and then recursing on each block. Since no perfect balanced partition is needed, GlobalMultisection has more freedom to move vertices, resulting in lower communication costs.
SharedMap [44] further decreases the communication cost. It introduces an adaptive imbalance strategy that rescales the imbalance parameter $\epsilon$ for each individual partitioning. It guarantees that the final $k$-way partition is $\epsilon$ balanced, and often gives each individual partitioning more space to work with. This results in an overall better mapping with lower communication costs. SharedMap uses KaFFPa [40] as one of its underlying graph partitioners, and it offers state-of-the-art solution qualities for process mapping.
The second approach to solve the GPMP, integrated mapping, combines the partitioning and mapping into one algorithm. It leverages the multilevel scheme and the refinement algorithms minimize $J(C,D,\Pi)$. Two prominent algorithm are IntegratedMapping [15] and Mt-KaHyPar-Steiner [24].
IntegratedMapping [15] uses matching based coarsening to create a smaller representation of the graph. The Global Path Algorithm [34], a $1/2$-approximation algorithm, is used to determine a matching of vertices. Instead of using the edge weights as a rating, a special rating function $\exp^{*}(\{u,v\})=\omega(\{u,v\})/c(u)c(v)$ from [26] is used. As initial partitioning, the hierarchical multisection approach is used. Refinement strategies include quotient graph refinement [40], $k$-way FM refinement [40], label propagation [38] and multi-try FM refinement [40]. All refinement methods use the gain of a vertex $v$ when moving it from its current block $\Pi(v)$ to a new block $b$. The gain is defined as follows:
$G_{b}(v)=\sum_{u\in N(v)}C_{v,u}\Big(D_{\Pi(v),\Pi(u)}-D_{b,\Pi(u)}\Big).$ (1)
Multiple methods are presented to query the distance $D_{i,j}$ between two PEs $i$ and $j$. The simplest one saves a matrix with $\mathcal{O}(k^{2})$ space and access each distance in $O(1)$. The other methods, have a considerable lower space complexity, but come with more computational load.
Mt-KaHyPar-Steiner [24] originally was developed to minimize wire lengths in VLSI designs. Logical units, hypergraph $\mathcal{H}$, are mapped to chip locations, edge-weighted graph $\mathcal{C}$, such that the total sum of used edge-weights, the implicit wire-length distance, is minimized. Replacing $\mathcal{H}$ with the task graph $G_{\mathcal{C}}$ and $\mathcal{C}$ with the hardware topology matrix $D$ leads to GPMP. Mt-KaHyPar-Steiner is integrated into the shared-memory parallel hypergraph partitioner Mt-KaHyPar [20]. The algorithm follows the multilevel (hyper-)graph partitioning approach. First the hypergraph is clustered, and these clustered are coarsed into a smaller graph. An initial $k$-way partition of the hypergraph is determined and then greedily mapped onto the chip locations. As local search refinement algorithms, label propagation [38], FM refinement [16] and flow-based refinement [40] are used.
III-C Kokkos
We utilize the Kokkos [14] library to implement our algorithms. Kokkos is a C++ library providing performance portability across many-core architectures, including GPUs. It abstracts two main concepts: (1) Parallel execution patterns and (2) Multidimensional arrays. Since multidimensional arrays are not used in this work, we focus on the parallel execution patterns, which are built around three fundamental primitives. (a) parallel_for executes $N$ independent work units in parallel. Each work unit operates on the same user defined function $f(i)$ with only the index $i\in[0,N)$ changing for each work unit. (b) parallel_reduce performs a reduction $R=\bigoplus_{i=0}^{N-1}f(i)$ over values returned by $f(i)$ using an associative operator $\oplus$. (c) parallel_scan computes an exclusive prefix sum over a range of values i.e. $y_{i}=\sum_{j=0}^{i-1}f(j)$. All work units share access to memory, and Kokkos provides atomic operations and deterministic reduction mechanisms to prevent race conditions and ensure correctness on many-core architectures.
III-D CSR Graph Format
A graph $G=(V,E)$ is typically stored in the Compressed Sparse Row (CSR) format (also referred to as Compressed Row Storage - CRS). The representation consists of an offset array $O$ (also called row-pointer) of size $|V|+1$ and two arrays of size $|E|$, one for the edge vertices $E_{v}$ and one for the edge weights $E_{w}$. The edges incident to vertex $v$ are stored in the contiguous range $[O[v],O[v+1]])$ of $E_{v}$ and $E_{w}$. Note that the graph is undirected and an edge $\{u,v\}$ is both stored for $u$ and $v$. This layout naturally supports parallelization across vertices, with each vertex processing its adjacency list independently.
IV GPU-Accelerated Process Mapping
The main contribution of this work is the development of two GPU-based approaches for solving the hierarchical process mapping problem. Given the close relationship between process mapping and graph partitioning, and the recent success of GPU-based partitioners in achieving faster runtimes, we aim to extend these advantages to process mapping.
Our first algorithm follows the hierarchical multisection approach. It recursively partitions the task graph $G_{\mathcal{C}}$ along the systems hierarchy: first across islands, then racks, and so forth. The final mapping of blocks, and thus vertices, to PEs is obtained by following the recursive calls. The algorithm is presented in Section IV-A.
The second algorithm integrates process mapping directly into a multilevel scheme and is inspired by the GPU-based graph partitioner Jet [18] (Section III-A). Its refinement phase has been adapted to employ the new gain equation (Equation 1). Further details are given in Section IV-B.
IV-A Hierarchical Multisection
The hierarchical multisection approach was first introduced by Kirchbach et al. [47]. It directly exploits the given hierarchy $H=a_{1}:\ldots:a_{\ell}$ of the supercomputer to partition the task graph $G_{\mathcal{C}}$. First, an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ is determined utilizing a graph partitioning algorithm. The resulting $a_{\ell}$ blocks, represent the largest hardware components of the supercomputer (e.g., the islands). Since the graph partitioner minimizes the edge-cut between the blocks, it directly corresponds to minimizing the communication cost across the islands. Then the algorithm recursively partitions each of the $a_{\ell}$ blocks into an $a_{\ell-1}$-way partition. These blocks represent the second-largest component of the system (e.g., racks within islands). Again, minimizing the edge-cut corresponds to minimizing the communication across the racks of each island. This recursive procedure is continued until $k$ blocks are obtained. The mapping of the $k$ blocks to the $k$ PEs can be extracted by following the recursive calls.
Schulz and Woydt [44] improved this method by introducing an adaptive imbalance parameter. Using the original imbalance parameter $\epsilon$ for each partitioning can overall result in an unbalanced $k$-way partition, i.e., some PEs may be overloaded. To avoid this problem, an adaptive imbalance parameter $\epsilon^{\prime}$ is determined for each partitioning. It depends on the original imbalance $\epsilon$, the overall graph weight $c(V)$, the weight of the current subgraph $c(V^{\prime})$, the desired number of blocks $k$, the number of total blocks the current subgraph will be partitioned into $k^{\prime}$ and the depth of the current partition in the hierarchy $d$. The adaptive imbalance is then determined by
$\epsilon^{\prime}=\Big((1+\epsilon)\frac{k^{\prime}c(V)}{kc(V^{\prime})}\Big)^{\tfrac{1}{d}}-1.$ (2)
Using the adaptive imbalance guarantees that the final partition is $\epsilon$-balanced. Experimental results in [44] demonstrate that the adaptive imbalance improves the overall mapping quality. In particular, the SharedMap [44] implementation currently achieves state-of-the-art solution quality for the HPMP.
To adopt the hierarchical multisection approach to GPUs, we must (1) employ a GPU-based graph partitioner and (2) construct the resulting subgraphs directly on the GPU.
For the partitioning step, we use Jet [18] which we already presented in Section III-A. Jet outperforms CPU-based graph partitioners in terms of speed by factor in the range from 2 to 20 times while delivering equal or better solution quality. Alternative GPU-based graph partitioners such as Sphynx [1] or G-kway [31] could also be used. As no comprehensive comparison between GPU partitioners exists, we adopt Jet for its thorough evaluation. However, any faster or higher-quality alternative would translate to our algorithm.
For the second step, we need to build the subgraphs of a graph $G=(V,E)$ given a partition $\Pi:[n]\mapsto[k]$. To circumvent the costly down- and upload between the CPU and GPU, the entire process is executed on the GPU. Our algorithm works in $k$ iterations, where each iteration builds one subgraph $G_{i}$. Each iteration consists of three phases. The first phase determines the number of vertices $n_{i}$, edges and the subgraphs weight using three parallel_reduce loops. A vertex $v$ belongs to the $i$th subgraph if $\Pi(v)=i$ and likewise an edge $\{u,v\}$ belongs to it, if $\Pi(u)=\Pi(v)=i$. The second phase determines a mapping between the vertices of $G_{i}$ to the vertices of $G_{\mathcal{C}}$. The vertices of $G_{i}$ need to be indexed from 0 to $n_{i}-1$, which is also incorporated into the mappings. The mappings are needed so we can propagate the assigned partition of vertices in $G_{i}$ back to the vertices of $G_{\mathcal{C}}$. This step is efficiently implemented via a parallel_scan over the vertex set. The prefix sum assigns each vertex a unique new vertex in the range [0,n1)[0,n-1). In the third phase, we construct the subgraph using a parallel_scan. The scan first computes the CSR offsets for each vertex and then writes the corresponding edges and weights. For each vertex $u$ only edges $\{u,v\}$ in the same subgraph are inserted, while the mapping functions translate between global and local vertex IDs. Subgraph creation usually takes less than 2% of the total runtime of the complete algorithm.
Algorithm 1 presents the pseudocode of the hierarchical multisection method. The function takes as input a (sub)graph $G_{L}$, the hierarchy $H$, the current hierarchy level $i$, an identifier list $I$, and a local mapping $\Pi_{L}$ that maps the vertices of $G_{L}$ to the vertices of $G_{\mathcal{C}}$. The identifier list $I$ uniquely identifies the position of the subgraph within the system hierarchy. Lines 242-4 handle the exit condition: when $i=0$, the subgraph represents one of the final $k$ blocks and all its vertices are assigned to exactly one of the $k$ PEs. The function $\textsc{calcID}(H,I)$ determines a unique ID for each identifier $I$ such that neighboring components receive consecutive IDs. The mapping $\Pi_{L}$ is used in Line 5 to translate the local vertices back to their original vertices in $G_{\mathcal{C}}$ and for these vertices the mapping is set to the ID.
If the subgraph $G_{L}$ resides on a higher level ($i>0$) then further partitioning is required (Lines 6126-12). First, The adaptive imbalance $\epsilon^{\prime}$ is determined and then Jet [18] is utilized to partition $G_{L}$ into $a_{i}$ subgraphs via the mapping $\Pi^{\prime}:V(G_{L})\mapsto[a_{i}]$. For each resulting subgraph $G^{\prime}_{L}$, we determine a local mapping $\Pi^{\prime}_{L}$ from its vertices to $G_{\mathcal{C}}$ (needed in Line 5). The identifier list $I$ is copied and extended by index $j$ to uniquely identifier the subgraph (needed in Line 3). The algorithm then recurses on each subgraph. Finally, the complete algorithm is invoked in Line 14 with the task graph $G_{\mathcal{C}}$, the hierarchy $H$, the highest hierarchy level $\ell$, an empty identifier list, and the identity mapping for $G_{\mathcal{C}}$.
Algorithm 1 Recursive Hierarchical Multisection
1:Graph $G$, Hierarchy $\smash{H=a_{1}:\ldots:a_{\ell}}$
[…]
3:  if $i=0$ then
4:   $\textsc{ID}\leftarrow\textsc{calcID}(H,I)$
5:   for $\smash{v\in V_{L}}$ do $\smash{\Pi(\Pi_{L}(v))\leftarrow\textsc{ID}}$
6:  $\smash{\epsilon^{\prime}\leftarrow\textsc{AdapImb}(\ldots)}$ $\triangleright$ uses Eq. 2
7:  $\smash{\Pi^{\prime}\leftarrow\textsc{Jet}(G_{L},a_{i},\epsilon^{\prime})}$
8:  $j\leftarrow 0$
9:  for $G^{\prime}_{L}\in\textsc{CreateSubgraphs}(G_{L},\Pi^{\prime})$ do
10:   $\Pi^{\prime}_{L}\leftarrow\textsc{MapToG}(G^{\prime}_{L},\Pi_{L})$
11:   $\smash{\textsc{HM}(G^{\prime}_{L},H,i-1,I+[j],\Pi^{\prime}_{L})}$
12:   $\smash{j\leftarrow j+1}$
13:end function
[…]
IV-B Integrated Mapping
The second main contribution of this work is an integrated mapping algorithm that incorporates the mapping directly into the graph partitioning. We follow the multilevel approach, but instead of minimizing the edge-cut we optimize $J(C,D,\Pi)$. The multilevel approach, coarsens the graph to a small representation, determines an initial partition and afterwards uncoarsens the graph and refines the mapping. We will detail our methodology for each phase in the next few paragraphs. We closely follow the matching, coarsening, refinement and rebalancing methods used by the GPU-based graph partitioner Jet [18] and adapt them to process mapping.
Edge-List Graph Format. Before presenting the multilevel process mapping scheme, we want to present our extended CSR format. As introduced in Section III-D, the CSR representation naturally supports parallelization over vertices, allowing each thread to process a vertex’s adjacency list independently. However, some computations benefit from direct parallelization over edges. In standard CSR, this requires nested parallel loops, first over vertices, then their neighbors, which leads to poor load balancing due to varying vertex degrees, resulting in idle threads on GPUs. To overcome this, we extend the CSR format with an additional array $E_{u}$ of size $|E|$ that explicitly stores the opposite endpoint of each edge. This edge-centric representation enables flat parallelism over all edges, improving load balance and GPU utilization. We refer to this edge-list representation as $\mathbb{E}$ throughout the rest of this work.
Matching. We utilize the two-hop matching presented by LaSalle et. al [29] but make a few changes. It starts with a heavy-edge matching algorithm. Instead of using $\omega(\{u,v\})$ as the weight of the edge, we use the rating function $\exp^{*2}(\{u,v\})=\frac{\omega(\{u,v\})^{2}}{c(u)c(v)}$ presented by Holtgrewe et al. [26] and also used in IntegratedMapping [15]. Their experiments showed that it lead to improved running times and solution quality in the context of graph partitioning. By looping over $\mathbb{E}$ in parallel, we determine for each vertex $v$ its preferred neighbor $p(v)=u$ as the one with the greatest rating. Adjacent vertices $u$ with $c(v)+c(u)>L_{\max}$ are excluded to prevent an imbalanced initial solution. Afterwards, we determine in parallel for each vertex $v$ if $p(p(v))=v$, i.e., vertex $v$ prefers $p(v)$ and $p(v)$ prefers $v$. These vertices are matched to each other. If the vertices have multiple neighbors with the same rating, then it is very unlikely that $v$ and $p(v)$ prefer each other, and we would end up with few matches. We therefore add a small amount of deterministic noise $\eta(\{u,v\})$ onto each rating. The noise breaks the equality and is small enough to not influence the order of two distinct ratings. The matching procedure can be repeated multiple rounds, excluding previously matched vertices. Afterwards, leaf, twin and relative matchings are also applied multiple times if necessary. We adopt the same hashing and matchmaker vertex strategy of Jet to improve performance. Instead of matching at least 75% of vertices, we reduce the percentage to 40%. While this leads to more levels in the multilevel framework, it also allows for more chances to refine the partition in the refinement phase. In our preliminary experiments using greater values lead to drastically worse solution quality.
Coarsening. Edge-contraction, to create the coarsened graph, is based on a fine-grained per-vertex hashing scheme from Jet [18]. The pseudocode can be seen in Algorithm 2. Let $G=(V,E)$ be the current graph and let $M:V\mapsto[n_{c}]$ map each vertex to its coarse counterpart. The value $n_{c}$ is the number of vertices in the coarse graph. The algorithm first computes an upper bound on the neighborhood size of each coarse vertex by summing the degrees of all vertices assigned to it (Lines 131-3). In Line 4 and 5 two hash arrays, $H_{v}$ and $H_{w}$, are allocated that will hold the edges and their weight. The offset array $O$ denotes the index range $O[v]$ to $O[v+1]$ where the edges of the coarse vertex $v$ are placed. Next all edges $(u,v,w)\in\mathbb{E}$ are processed in parallel. If both endpoints contract to the same vertex, the edge becomes a self loop and is discarded (Line 8). Otherwise we need to insert $(M(v),w)$ into the adjacency of $M(u)$ in the interval $\mathcal{I}$. Note that adding $(M(u),w)$ into the adjacency of $M(v)$ is handled by the work unit with reverse edge $(v,u,w)\in\mathbb{E}$. Insertion is done in Line 10, which we will explain in the next paragraph. After the parallel loop ends it remains to extract the actual graph from the arrays $H_{v}$ and $H_{w}$. This boils down to determining the true neighborhood sizes of each vertex and extracting the edges by skipping all Null entries in $H_{v}$.
The most important part of our fine grained edge-contraction is the InsertOrLookup function in Line 10. To handle concurrent insertions without locks we utilize atomic compare-and-swap (CAS) operations. A CAS operation compares the value at $H_{v}[j]$ with Null and only if so swaps $M(v)$ in its place. Regardless if the check succeeded the previous value is returned. This means, if a thread executes a CAS operation on $H_{v}[j]$ and Null or $M(v)$ is returned, a slot for $M(v)$ was found. The weight $w$ will be atomically added to $H_{w}[j]$. Otherwise the thread will continue to cycle through the interval $\mathcal{I}$. Since we overestimated the neighborhood size of $M(u)$ at least one open slot will be available. To reduce atomic contention on $H_{v}[j]$, each thread determines $j\in\mathcal{I}$ by hashing $M(v)$ and then cycling through the range if necessary.
Algorithm 2 GPU-based Coarsening
1:Graph $G$, Mapping $M$, # Coarse Vertices $n_{c}$
[…]
3:for $v\in V$ do in parallel
4:  $\smash{B[M(v)]\leftarrow B[M(v)]+|N(v)|}$
[…]
8:for $(u,v,w)\in\mathbb{E}$ do in parallel
9:  if $M(u)=M(v)$ then skip $\triangleright$ Self-loop discarded
10:  $\smash{\mathcal{I}\leftarrow[O[M(u)],O[M(u)+1])}$
11:  $\smash{\textsc{InsertOrLookup}(H_{v},H_{w},\mathcal{I},M(v))}$
12:return $\smash{\textsc{ExtractCSR}(B,H_{v},H_{w})}$
Initial Partitioning. Initial partitioning is performed once the last coarsened graph has fewer than $128k$ vertices. Choosing a smaller threshold often resulted in an imbalanced initial partitions from which the algorithm failed to recover. We apply the hierarchical multisection approach explained in Section IV-A to acquire an initial mapping, but execute it entirely on the CPU. At this stage, the graph is sufficiently small that GPU execution offers no runtime advantage. Moreover, even the GPU-based partitioner Jet [18] would internally quickly fall back to the CPU-based partitioner Metis [27] when presented with such small graphs. To execute hierarchical multisection, we utilize the CPU-based partitioner KaFFPa [40] in its Fast configuration. Although the Strong configuration yields higher-quality mappings, its significantly increased runtime would dominate the total execution time, rendering the overall algorithm non-competitive.
Refinement. We adapt the unconstrained label propagation of Jet [18] to use $J(C,D,\Pi)$ as the objective function. The corresponding pseudocode is provided in Algorithm 3. The refinement algorithm begins by determining the best target block $\Pi^{\prime}(v)$ for each vertex $v$ by maximizing the gain $G_{b}(v)$ (see Equation 1) across all neighboring blocks (Lines 353-5).
Algorithm 3 Unconstrained Label Propagation
1:Graph $\smash{G}$, Mapping $\Pi$, Is-Locked array $L$
[…]
4:for $v\in V$ do in parallel
5:  if $\smash{L[v]=1}$ then skip $\triangleright$ Vertex locked
6:  $\smash{B\leftarrow\{\Pi(u)\mid u\in N(v)\}}$
7:  $\smash{\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)}$
8:  if $\smash{G_{\Pi^{\prime}(v)}(v)\geq 0}$ then $\smash{\mathcal{X}[v]=1}$
[…]
10:for $v\in V$ do in parallel
11:  if $\smash{\mathcal{X}[v]=1\text{ and }\smash{\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0}}$ then $\smash{\mathcal{M}[v]=1}$
[…]
13:return $\smash{\mathcal{M},\Pi^{\prime},L}$
In Jet, a filter is applied that retains only vertex moves satisfying $G_{b}(v)\geq 0\lor-G_{b}(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. This criterion is suitable when minimizing the edge-cut, as both $G_{b}(v)$ and $\operatorname{conn}(v,\Pi(v))$ depend solely on edge weights and are thus directly comparable. In contrast, for process mapping, $G_{b}(v)$ contains the additional communication factors and is therefore often much smaller than $\operatorname{conn}(v,\Pi(v))$ and varies significantly, depending on $v$’s connections. This renders the original condition ineffective. We therefore only allow non-negative moves $G_{b}(v)\geq 0$ to pass the filter, as it performed best in our preliminary experiments. A more sophisticated filtering criterion may further improve solution quality and remains a topic for future work.
In Line 7, vertices passing this filter are inserted into the set $\mathcal{X}$, which is implemented as a binary array marking selected vertices. The set is implicitly ordered by $\operatorname{ord}$, defined as
$\operatorname{ord}(u)\begin{cases}<\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)>G(v),\\ <\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)=G(v)\land u<v,\\ >\operatorname{ord}(v)&\text{otherwise}.\end{cases}$
Next, all vertex moves in $\mathcal{X}$ are reevaluated under an approximated future mapping state. If a vertex $v$ has a neighbor $u$ scheduled to move to block $\Pi^{\prime}(u)$ with $\operatorname{ord}(u)<\operatorname{ord}(v)$, then $u$ is assumed to have already moved when recomputing the gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$. We mark each vertex $v\in\mathcal{X}$ in a To-Move array $\mathcal{M}$ if it has non-negative gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ (Lines 9109-10). To prevent oscillation of vertices, the To-Move array serves as the locked array $L$ for the next iteration (Line $11$). The function returns the To-Move array $\mathcal{M}$ the destination mapping $\Pi^{\prime}$ and the Locked array $L$.
Rebalancing. Our rebalancing is also adapted from Jet [18]. Jet splits its rebalancing into a weak and strong algorithm. The weak rebalancing keeps the loss small, however, it might need multiple iterations to achieve a balanced partition. The strong rebalancing instead often achieves a balanced partition in one iteration, but has a greater loss.
Weak rebalancing starts by determining for each vertex $v$ in an overloaded block, the move with maximum gain (or equally minimum loss) $G_{b}(v)$ into a neighboring block $b$ that has a weight smaller than the threshold $\sigma\leq L_{\max}$. The threshold $\sigma$ is chosen such that there is a small dead-zone between $\sigma$ and $L_{\max}$, so $b$ can take multiple vertices without instantly being overloaded. In our implementation we start with $\sigma=L_{\max}-0.065L_{\max}$ and gradually lower it throughout the multilevel scheme until at the lowest level we use $\sigma=L_{\max}-0.005L_{\max}$. We lower it, since the vertex weights are also gradually getting smaller during uncoarsening. Therefore, we can operate on a more fine level and more blocks become eligible. If for a vertex no such neighboring block exists, a random block that fits the threshold is chosen.
Ideally, each block would store all its outgoing vertex moves decreasingly in a list and then a prefix of the list could be chosen, such that the block becomes balanced. This would grant the smallest possible loss based on the current partition. However, since sorting can be expensive on a GPU, we utilize an approximation of the sorted list similar to Jet. Each vertex move is assigned a slot in the list $0,-1,\ldots,-10,-20,\ldots,-100,-200,\ldots,-1000,\operatorname{X}]$. The slot ++ holds all positive gain moves, the slot $\operatorname{X}$ contains all very large negative gain moves and all other slots hold moves with gain $L[i]\geq G_{b}(v)>L[i+1]$. Note that at first we increase the range by 1, then by 10 and in the end by 100. This gives us an increased resolution for moves with a comparatively small loss, which minimizes the loss when only a few vertices need to be moved. Note that Jet also adopts the ++ and 0 slot and for any negative moves they use $\log_{2}(-G_{b}(v))$ to assign a vertex a slot. The approximated list is build in four phases, first by counting how many vertices fall into each slot, allocating enough memory, building a exclusive prefix-sum over the sizes and in the last phase inserting the vertices into their slots. To reduce atomic contention we utilize $\rho=2$ mini buckets in each slot and a vertex $v$ is assigned its mini bucket by $v\mod\rho$. As the last step of weak rebalancing a list prefix is chosen for each block $b$ and all vertices in it are moved. The prefix moves at least $c(b)-L_{\max}$ weight, so that the block $b$ becomes balanced.
Since the lists are approximately sorted by their gain, the prefixes have a small combined loss. However, there is no safeguard against the destination blocks becoming overloaded. While the threshold $\sigma$ helps in preventing this, it does not guarantee it. Therefore, the destination blocks might become overbalanced and another iteration is required.
Strong rebalancing operates in a similar manner to weak rebalancing but reverses the assignment perspective. As before, each vertex $v$ in an overloaded block $b$ identifies an eligible neighboring target block $b^{\prime}$ that minimizes the loss of moving $v$. However, instead of storing the vertex moves with the source block $b$, they are stored in the lists of the corresponding target block $b^{\prime}$. Each target block then selects a prefix of moves from its own list, such that it stays balanced after accepting them. After execution, the target blocks are guaranteed to stay balanced, while the source blocks lose weight and are also likely to be balanced. Consequently, strong rebalancing can restore balance in just one iteration. However, target blocks will often move more vertices than necessary, resulting in a higher overall loss compared to weak rebalancing.
Algorithm 4 Overall Refinement Algorithm
1:Graph $G=(V,E)$, Mapping $\Pi$
[…]
5:while $i<i^{\max}$ do
6:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
7:   $\mathcal{M},$ $\Pi^{\prime\prime}$, $L$ $\leftarrow\operatorname{UnconLabelProp}(G,\Pi^{\prime},L)$
8:   $i_{w}\leftarrow 0$
9:  else
10:   $L\leftarrow\textsc{Zero}(|V|)$ $\triangleright$ Reset Locks
11:   if $i_{w}<i^{\max}_{w}$ then
12:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{weakReb}(G,\Pi^{\prime})$
13:     $i_{w}\leftarrow i_{w}+1$
14:   else
15:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{strongReb}(G,\Pi^{\prime})$
16:     $i_{w}\leftarrow 0$
17:  $\Pi^{\prime}\leftarrow\operatorname{move}(\mathcal{M},\Pi^{\prime\prime})$
18:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
19:   if $J(C,D,\Pi^{\prime})<J(C,D,\Pi)$ then $\Pi\leftarrow\Pi^{\prime}$
20:   if $J(C,D,\Pi^{\prime})<\phi J(C,D,\Pi)$ then $i\leftarrow 0$
21:  else
22:   if $\operatorname{maxImb}(G,\Pi^{\prime})<\operatorname{maxImb}(G,\Pi)$ then
23:     $\Pi\leftarrow\Pi^{\prime},\quad i\leftarrow 0$
Overall Refinement Algorithm. The complete refinement procedure is summarized in Algorithm 4. The algorithm iteratively alternates between unconstrained label propagation and rebalancing (depending on the current imbalance) until the maximum number of iterations $i^{\max}$ is reached. A copy of the mapping $\Pi^{\prime}$ is used and the actual mapping $\Pi$ is only updated if an improvement mapping has been found. The parameter $i^{\max}_{w}$ controls how many consecutive iterations of weak rebalancing are applied before a single iteration of strong rebalancing is triggered.
In each iteration, if the current partition $\Pi^{\prime}$ satisfies the imbalance constraint (Line 5), the algorithm performs unconstrained label propagation to reduce the communication cost (Line 6). Otherwise, the algorithm invokes weak or strong rebalancing to improve the balance (Lines 8158-15). After each iteration, the partition $\Pi^{\prime}$ is updated (Line 16), and the best feasible solution encountered so far is saved.
The process terminates when no significant improvement in either objective quality or balance has been achieved within the iteration limits. The hyperparameter $\phi=0.999$ specified the minimum relative improvement of $J(C,D,\Pi)$ that has to be achieved in each iteration. We set $i^{\max}_{w}=2$ by default, but increase it to $10$ in the final refinement stage of the multilevel pipeline to ensure minimal loss while achieving balance. The maximum number of iterations $i^{\max}$ is set to $12+\operatorname{level}$, where $\operatorname{level}$ denotes the current level in the coarsening hierarchy (the original graph at level 0, the coarser representation at higher levels). At coarser levels, graphs are smaller, allowing more iterations, but at finer levels, graphs are larger, so each iteration is more computationally expensive.
Additional Data Structure. Determining the gain of moving a vertex to its neighboring blocks is essential for both label propagation and rebalancing. While this information can be obtained by iterating over the neighbors of a vertex, it can be inefficient as a vertex with 100 neighbors might only be connected to two blocks. We employ an additional data structure [18] that enables rapid access to the neighboring blocks. Conceptually, it is similar to the CSR representation of a graph, however, instead of storing edges, it stores the neighboring blocks and their the aggregated weight of edges. For each vertex $v$, we initially allocate memory for $|\{\Pi(u)\mid u\in N(v)\}|+2$ entries. The additional capacity accounts for potential future vertex moves. Note that this is less memory than Jet, which allocates $\min\{k,|N(v)|\}$ per vertex. Whenever a vertex changes its block assignment, we attempt to update its connectivity within the existing memory. We update the connectivity of $u$ to its neighbors $v$ as follows. For each $(u,v,w)\in\mathbb{E}$, we subtract $w$ from $v$’s connection to $\Pi(u)$. Should the resulting weight drop to 0, then the slot is freed. Next, we add the connectivity to $u$’s new block $\Pi^{\prime}(u)$ in $v$’s connectivity, if $\Pi^{\prime}(u)$ does not already exist. If this is not possible, because all slots are already occupied, we mark in a bit array, that $\Pi^{\prime}(u)$ could not be inserted into $v$’s adjacency. As a last step we add the weights $w$ to the slot $\Pi^{\prime}(u)$. All these steps utilize a parallel_for loop over the edge list $\mathbb{E}$. If any vertex cannot be updated because not enough slots are available, we trigger a full rebuild of the data structure. Any vertex that exceeded its capacity receives $|\{\Pi(u)\mid u\in N(v)\}|+2$ slots, while all other vertices retain their slots.
V Experimental Evaluation
All of our algorithms111github.com/HenningWoydt/GPU-HeiProMap are implemented in C++ 17 and utilize the Kokkos library [14] version 4.7.0 for the GPU code. It compiles the host side code with GCC 11.4 and the GPU code with Nvidia Cuda compiler 12.6. We use KaFFPa [40] from the KaHIP library version 3.19.0, which is also implemented in C++. We compile with the highest optimization (-O3).
System. The experiments are conducted on a machine that has an Intel® Xeon® w5-3435X, that runs on a max frequency of 3.13.1Ghz. The system features 128 GB of main memory. The used GPU is a NVIDIA® GeForce RTX™ 4090 with 24GB of memory. It has 16.384 Cuda cores with a boost frequency of 2.52.5Ghz. The system runs Ubuntu 22.04.4.
Benchmark Instances. To make our work more comparable, we choose the same benchmark instances as in [44, 15, 24, 43, 47]. They can be seen in Table I. The small instances come from the SuiteSparse Matrix Collection [11]. These represent task graphs from real application. Six graphs were taken from Walshaws’s benchmark archive [45], which holds instances for benchmarking graph partitioning algorithms. The graphs af_shell9, thermal2, nlr are from the matrix and numerical section of the 10th DIMACS Implementation Challenge [5]. The graphs deu and eur are large road networks of Germany and Europe [12]. Graphs rgg23 and rgg24 are random geometric graphs with $2^{23}$ and $2^{24}$ vertices respectively. Each vertex represents a random point in the unit square, and if their distance is below than $0.55\sqrt{\ln n/n}$ they are connected via an edge. Graphs del23 and del24 also have $2^{23}$ and $2^{24}$ vertices, with each vertex representing a point in the unit square. The edges are connected such that the graph is Delaunay triangulation.
TABLE I: Benchmark instance properties.
SuiteSparse Other
Graph $|V|$ $|E|$ Graph $|V|$ $|E|$
cop20k_A 99 843 1 262 244 af_shell9 $\approx$504K $\approx$8.5M
2cubes_sphere 101 492 772 886 thermal2 $\approx$1.2M $\approx$3.7M
thermomech_TC 102 158 304 700 nlr $\approx$4.2M $\approx$12.5M
cfd2 123 440 1 482 229 deu $\approx$4.4M $\approx$5.5M
boneS01 127 224 3 293 964 del23 $\approx$8.4M $\approx$25.2M
Dubcova3 146 689 1 744 980 rgg23 $\approx$8.4M $\approx$63.5M
bmwcra_1 148 770 5 247 616 del24 $\approx$16.7M $\approx$50.3M
G2_circuit 150 102 288 286 rgg24 $\approx$16.7M $\approx$132.6M
shipsec5 179 860 4 966 618 eur $\approx$18.0M $\approx$22.2M
cont-300 180 895 448 799
Walshaw
598a 110 971 741 934 wave 156 317 1 059 331
fe_ocean 143 437 409 593 m14b 214 765 1 679 018
144 144 649 1 074 393 auto 448 695 3 314 611
(a) Solution quality
(b) Speedup
Algorithm Phase Small Large GPU-HM I/O 24.9% 47.9% Jet 75.1% 52.1% GPU-HM-ultra I/O 4.8% 13.7% Jet 95.2% 86.3% GPU-IM I/O 29.3% 74.9% Heavy-Edge Matching 0.4% 1.1% Edge-Contraction 0.4% 4.7% KaFFPa 64.8% 6.2% Label Prop. + Bal. 5.0% 12.9% Uncoarsening 0.1% 0.2%
(c) Runtime breakdown
Figure 1: Comparison of solution quality (a) and speedup (b) of our proposed algorithms. Table (c) shows the runtime breakdown per phase.
Experimental Setup. As the hierarchy we choose $H=4:8:\{1,\ldots,6\}$ and as the distance we choose $D=1:10:100$. The imbalance parameter is set to 3% ($\epsilon=0.03$). Each algorithm is run five times with different seeds and the running time and communication cost $J(C,D,\Pi)$ are averaged across the seeds.
Performance Profiles. Performance Profiles [13] enable us to analyze the quality of multiple algorithms. Let $\mathcal{A}$ be the set of all algorithms and let $\mathcal{I}$ be the set of all instances. The quality of algorithm $A\in\mathcal{A}$ on instance $I\in\mathcal{I}$ is $q_{A}(I)$ and $\text{Best}(I)=\min_{A\in\mathcal{A}}q_{A}(I)$ denotes the best quality across all algorithms on instance $I$. For each algorithm $A$, a performance plot shows the fraction of instances, the $y$-axis, for which $q_{A}(I)\leq\tau\text{Best}(I)$, where $\tau\geq 1$ is on the $x$-axis. For $\tau=1$ the plot shows the fraction of instances that each algorithm solved with the best solution. Algorithms that perform well have a high fraction of instances solved with minimal $\tau$ values, they are towards the top-left corner. Sub-optimal algorithms will have small fractions at minimal $\tau$ values, and only achieve higher fractions with greater $\tau$ values.
V-A Results
We will first analyze the results of our GPU-based hierarchical multisection (GPU-HM) and integrated mapping (GPU-IM) approach. Note that Jet [18], which we utilize in GPU-HM, has an ultra configuration which executes 18 refinement iterations instead of only one, to improve solution quality. We abbreviated our hierarchical multisection approach that uses Jet-ultra as GPU-HM-ultra. In the end, we will compare our methods against the established algorithms SharedMap [44] and IntegratedMapping [15], as they are CPU-based equivalent algorithms to hierarchical multisection and integrated mapping.
Comparison. Figure 1 shows the performance profile (a) and the speedup (b) of our algorithms across all instances. GPU-HM-ultra serves as a baseline to measure the speedup as it achieves the lowest communication cost. The default GPU-HM approach reaches a maximum speedup of $6.4$ with a geometric mean speedup of $4.4$. However, the speedup comes with an increase in communication cost. While the ultra variant finds the best solution on 94.794.7% of instances, GPU-HM does so on only 5.35.3%. On average, the communication cost of GPU-HM-ultra is just 0.10.1% above the best solution, whereas GPU-HM is 9.39.3% worse. GPU-IM is the fastest of our algorithms with a maximum speedup of $16$ and a geometric mean speedup of $5.6$. However, it is also the worst in terms of communication cost, as it did not achieve the lowest communication cost on any instance and has on average 17.417.4% higher communication cost than the best solution.
Table 1c reports the percentage of runtime each algorithm spends in its major phases. Results are grouped by graph size: small graphs with fewer than one million vertices and large graphs with more than one million vertices. For the hierarchical multisection approach, most runtime is spent in the partitioning, which gets executed by Jet [18]. Only for large graphs does I/O become significant, accounting for 47.947.9% of the runtime in the case of GPU-HM. The I/O phase includes reading the graph from disk, constructing its data structures, transferring it to the GPU and writing the final partition to disk. Among these steps, loading and constructing the graph dominate the I/O runtime. Overall, reducing the runtime of Jet or employing a faster GPU-based partitioner appears to be the most promising direction for improving performance further. The runtime distribution for GPU-IM follows a similar pattern. The greatest runtime percentages are in the phases that do not utilize the GPU, namely I/O and utilizing KaFFPa [40] for the initial partitioning. For smaller graphs, one could coarsen the graph further to partition a smaller graph, which would reduce the required time. However, this often leads to imbalanced solutions and worse communication cost. As later results will show, GPU-IM already lags behind other CPU-based algorithms in terms of solution quality. We therefore utilize GPU-IM as a very fast but low-accuracy algorithm. For larger graphs, the integrated mapping approach spends most of its runtime in the I/O phase.
Figure 2: Comparison of speedup (left) and performance profile of achieved communication costs (right) of our proposed algorithms GPU-HM-ultra and GPU-IM to CPU-based algorithms SharedMap and IntegratedMapping.
Comparison to CPU-based Algorithms. We compare our algorithms to the CPU-based solvers SharedMap [44] and IntegratedMapping [15], since they performed well in the experimental evaluation of [44]. SharedMap employs the hierarchical multisection approach, just like GPU-HM. For partitioning, it internally uses KaFFPa [40] in either the Fast or Strong configuration, which trade speed for solution quality. We abbreviate the variants by SharedMap-F and SharedMap-S. IntegratedMapping, in contrast, minimizes $J(C,D,\Pi)$ via the multilevel scheme, making it conceptually similar to GPU-IM. It also provides a Fast and Strong configuration, which we denote by IntMap-F and IntMap-S. For comparison, we focus on GPU-HM-ultra, which achieves the highest solution quality, and GPU-IM, which provides the lowest runtime. GPU-HM is omitted for better visual clarity, as its performance lies between the two.
Figure 2 shows the speedup over SharedMap-S (left) and the performance profile (right). Among all algorithms, SharedMap-S achieves the highest solution quality, obtaining the best solution on 125 instances (83.383.3%). However, this comes at the cost of runtime, as SharedMap-S is also by far the slowest algorithm. GPU-HM-ultra finds the best solution on the remaining 25 instances (16.716.7%). All other algorithms are not able to determine a best solution on any instance. Ranked by average additional cost over the best solution, the algorithms are ordered as follows: SharedMap-S (0.020.02%), GPU-HM-ultra (11.511.5%), IntMap-S (14.514.5%), IntMap-F (21.121.1%), GPU-IM (30.730.7%) and SharedMap-F (31.731.7%). GPU-IM and SharedMap-F are at most $2.42\times$ and $2.66\times$ worse than the best-known solution, respectively.
In terms of runtime, GPU-IM is the fastest algorithm overall, achieving a maximum speedup of $598.5$ and a geometric mean speedup of $77.7$ compared to SharedMap-S. While GPU-HM-ultra achieves a maximum speedup of $377$, its geometric mean speedup is only $13.8$. On many instances, it is slower than IntMap-F/S and SharedMap-F, largely due to the partitioning phase being the bottleneck (see Table 1c). The remaining mean speedups are $28.4$ for SharedMap-F, $23.8$ for IntMap-F, and $10.4$ for IntMap-S.
Our algorithms have two clear equal CPU-based counterparts. GPU-HM-ultra slightly outperforms IntMap-S in terms of solution quality. Its geometric mean speedup of $13.8$ also exceeds the $10.4$ of IntMap-S, although it is not faster across all instances. Specifically, IntMap-S is faster on 7171% of small instances (speedup of $1.4$), whereas GPU-HM-ultra dominates on 9696% of the large instances (speedup of $5.6$). Similarly, GPU-IM is comparable to SharedMap-F in terms of solution quality. On small graphs, it achieves on average 2.62.6% lower communication costs, while on larger graphs it performs about 2.32.3% worse. The geometric mean speedup of GPU-IM over SharedMap-F is $1.4$ and $10.6$ for small and large graphs respectively.
In summary, GPU-HM-ultra outperforms the CPU-based algorithm IntMap-S in both solution quality and running time. However, it cannot reach the state-of-the-art solution quality of SharedMap-S. Nevertheless, GPU-HM-ultra is on average an order of magnitude faster. GPU-IM can be almost up to $600$ times faster than SharedMap-S, but its solution quality is on average 30.730.7% higher than the best-known solution.
Comparison to GPU-based Algorithms. While no other GPU-based algorithms currently exist for solving the GPMP, we can compare our algorithms to Jet [18]. Note that Jet solves the graph partitioning problem, which corresponds to using a distance vector of $D=1:\ldots:1$. Since we test the distance $D=1:10:100$, Jet’s resulting partitions are unfit and the corresponding communication costs are worse than those of our algorithms. In terms of runtime, however, Jet is comparable to GPU-IM as they both follow the multilevel scheme. We base our comparison on the large graph instances ($\geq 1$ million vertices). Using GPU-IM as the baseline for runtime and communication costs, we observe that Jet and Jet-ultra incur additional communication costs of 42.142.1% and 38.838.8% on average respectively. In terms of runtime, Jet-ultra is comparable to GPU-IM, with a geometric mean speedup of $1.03$, whereas Jet achieves an average speedup of $1.33$. We attribute this performance gap primarily to the 4040% matching requirement of our algorithm, compared to the 7575% in Jet. This allows Jet to produce smaller coarse graphs in fewer levels, and thereby reduce runtime. While we could tune our constants for more speed, we observed that it negatively impacted solution quality. Our algorithms GPU-HM and GPU-HM-ultra further improve solution quality but are worse in runtime than GPU-IM (see Figure 1), which translates to this GPU comparison.
VI Conclusion
Process mapping assigns vertices of a task graph to the processing elements (PEs) of a supercomputer to balance workload and minimize communication costs. Motivated by the success of GPU-based graph partitioners, we proposed two GPU-accelerated algorithms for this problem. The first, GPU-HM, applies hierarchical multisection [47] using the GPU-based partitioner Jet [18]. The second, GPU-IM, integrates process mapping directly into a multilevel framework, adapting Jet’s label propagation and rebalancing. GPU-HM-ultra achieves up to $300\times$ speedups and is on average a magnitude faster than the state-of-the-art SharedMap [44], with only 11.511.5% higher communication costs and therefore outperforming IntMap-S [15]. GPU-IM delivers extreme speedups (mean $77.7\times$, max $598.5\times$) at lower solution quality, however it still outperforms the Fast configuration of SharedMap. Future work will focus on improving solution quality without sacrificing performance e.g., enhancing label propagation filters, refining coarsening and initial partitioning, and exploring more efficient GPU-based partitioners.

### v2

@@
-GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1
+GPU-Accelerated Algorithms for Process Mapping
 Abstract
@@
 Vital phases like coarsening and refinement are accelerated by exploiting the parallelism of GPUs.
-In our experiments, both methods achieve speedups exceeding 300 when compared to state-of-the-art CPU-based algorithms.
-The first algorithm has, on average, about 10 percent greater communication costs and thus remains competitive to CPU algorithms.
-The second approach is much faster, with a geometric mean speedup of 77.6 and peak speedup of 598 at the cost of lower solution quality.
+The first algorithm has, on average, about 12 percent higher communication costs than the state-of-the-art solver and thus remains competitive with it.
+However, in terms of speed, it vastly outperforms the competitor with a geometric mean speedup of 22 times and a maximum speedup of 934 times.
+The second approach is even faster, with a geometric mean speedup of 1454 times and a peak speedup of 12376 times.
+Compared to other algorithms that prioritize speed over solution quality, this approach has the same quality but much greater speedups.
 To our knowledge, these are the first GPU-based algorithms for process mapping.
-I Introduction
+keywords:
+GPU, Process Mapping, Graph Partitioning
+1 Introduction
 Parallel high-performance applications rely on the efficient execution of large numbers of tasks across massively parallel architectures.
@@
 In parallel applications, the individual tasks usually communicate with each other, i.e., they send and receive data to and from each other.
-The cost of this communication heavily depends on the location of the tasks in hardware topology.
-Modern supercomputers are typically organized hierarchically, for example into processors, nodes, racks and islands.
+The cost of this communication heavily depends on the location of the tasks in the hardware topology.
+Modern supercomputers are typically organized hierarchically, for example into processors, nodes, racks, and islands.
 The communication cost between two PEs, therefore, depends on their relative locations within this hierarchy.
-Two PEs on the same processor communicate data much more quickly than two PEs which are located on different islands.
+Two PEs on the same processor communicate data much more quickly than two PEs that are located on different islands.
 Thus, in addition to balancing the workload, it is advantageous to minimize the total communication cost by placing highly interactive tasks close to each other.
@@
 Optimally solving the GPMP is $\mathrm{NP}$-hard [40] as it involves solving the well-known quadratic assignment problem (QAP) [10].
-The QAP asks for a set of $n$ facilities and $n$ locations to minimize the total cost of assigning each facility to one location, where the cost depends on both the flow between facilities and the distance between their locations.
+The QAP considers a set of $n$ facilities and $n$ locations.
+The objective is to assign each facility to exactly one location in a way that minimizes the total cost.
+This cost is determined by the flow between pairs of facilities and the distance between the locations to which they are assigned.
+In other words, facilities with higher interaction should ideally be placed closer together to reduce the overall cost.
 Since the QAP is strongly $\mathrm{NP}$-hard, i.e., no constant factor approximation exists unless $\mathrm{P}=\mathrm{NP}$, its hardness transfers to the GPMP.
 The hardness is also reflected in practice as no meaningful instance with $n>20$ can be solved in a reasonable time [10].
-Moreover, the same hardness result arises from graph partitioning (GP), where the goal is to partition a graph while minimizing the number of edges between them.
-Graph partitioning is $\mathrm{NP}$-complete and also admits no constant-factor approximation unless $\mathrm{P}=\mathrm{NP}$.
+Moreover, the same hardness result arises from graph partitioning (GP).
+GP considers a graph $G=(V,E)$ and asks for a partition of the vertex set into $k$ pairwise disjoint sets of equal size, such that the number of edges crossing between the different subsets is minimized.
+Graph partitioning is $\mathrm{NP}$-complete [18] and also admits no constant-factor approximation unless $\mathrm{P}=\mathrm{NP}$ [4].
 The GPMP is commonly solved using either the two-phase approach or via integrated mapping.
 In the two-phase approach, the task set is first partitioned into as many (almost) equally sized sets as there are PEs in the supercomputer.
-While partitioning the task set, the necessary communication between the resulting sets is minimized.
+While partitioning the task set, the necessary volume of communication across the different resulting sets is minimized.
 In the second phase, the subsets are mapped to the PEs such that the total cost of communication is minimized, taking the distances between the PEs into account.
-The first phase is an instance of GP, while second phase is a QAP instance.
-The two-phase approach is often easier to realize as it can fall back to two well-studied problems, i.e., graph partitioning the QAP.
-The algorithm SharedMap [44] currently offers state-of-the-art solution quality by utilizing hierarchical multisection, a special variant of the two-phase approach.
-The integrated mapping approach combines both phases into a single process by incorporating the mapping objective directly into the graph partitioning pipeline.
-Rather than minimizing the number of edges crossing between the subsets, as in classical graph partitioning, it directly minimizes the total cost of communication.
-Typically, this is achieved using the modern multilevel framework, which involves coarsening the input graph to a smaller representation, computing an initial mapping, and then refining it during uncoarsening.
-The approach is more complex to realize as the whole graph partitioning pipeline has to be adopted to the new objective function.
-However, this also enables function specific optimization, which result in an overall faster execution time.
-As a result, integrated mapping often achieves faster runtimes compared to the two-phase approach.
-Recently, Graphic Processing Units (GPUs) have been utilized to tackle graph partitioning.
-GPUs provide massive parallelism that can significantly reduce runtime, when exploited effectively.
-However, leveraging thousands of GPU cores is challenging due to the irregularity of graphs and lack of straightforward parallelization strategies.
+While the first phase is an instance of graph partitioning, the second phase is an instance of the quadratic assignment problem.
+This makes the approach easier to realize as one can fall back to two well-studied problems.
+The algorithm SharedMap [45] currently offers state-of-the-art solution quality by utilizing hierarchical multisection.
+This special variant of the two-phase approach directly exploits the hardware topology of the supercomputer.
+The integrated mapping approach combines both phases into a single process.
+The mapping objective of the second phase is directly integrated into the partitioning of the first phase.
+Rather than minimizing the communication volume between the subsets, it directly minimizes the total cost of communication.
+Typically, this is achieved using the modern multilevel graph partitioning framework.
+The input graph is first coarsened to a smaller representation and then an initial partition is determined.
+During uncoarsening, the partition is refined using local search algorithms.
+The approach is more complex as the whole graph partitioning pipeline has to be adopted to the new objective function.
+However, this also enables function-specific optimization resulting in an overall faster running time.
+Integrated mapping algorithms, like IntMap [16], often achieve faster running times compared to the two-phase approach.
+Recently, Graphics Processing Units (GPUs) have been utilized to tackle graph partitioning.
+GPUs provide many more cores than CPUs and, when exploited efficiently, can significantly reduce the running time.
+However, effectively leveraging this level of parallelism for graph algorithms is challenging.
+Graph computations typically involve irregular memory access patterns, uneven workloads across vertices, and complex data dependencies.
 Despite this, recent GPU-based algorithms like Jet [19] and G-kway [32] have demonstrated promising results.
-They deliver solution qualities comparable to the serial CPU-based partitioner Metis [27], while being up to 20 times faster.
-Since both the two-phase and integrated mapping approach rely on graph partitioning, these results should transfer to GPMP.
-[…]
-II Preliminaries
-II-A Concepts and Notations
+They deliver solution qualities comparable to the serial CPU-based graph partitioner Metis [28] while being up to 20 times faster.
+Both the two-phase and integrated mapping approach rely on graph partitioning, so these results should transfer to GPMP.
+Our Contribution.
+We present two GPU-based algorithms to solve the process mapping problem.
+The first one is based on the two-phase approach and utilizes hierarchical multisection.
+In contrast, the second algorithm integrates the mapping into the graph partitioning pipeline.
+Both significantly outperform state-of-the-art CPU solvers in runtime, achieving speedups of more than $900\times$, while remaining competitive with their solution quality.
+2 Concepts and Notations
 Let $G=(V,E)$ with $n=|V|$ and $m=|E|$ be an undirected graph, and let $c:V\mapsto\mathbb{R}$ and $\omega:E\mapsto\mathbb{R}$ be vertex and edge weights.
-Their extension to sets is $c(V^{\prime})=\sum_{v\in V^{\prime}}c(v)$ and $\omega(E^{\prime})=\sum_{e\in E^{\prime}}\omega(e)$ for vertex sets $V^{\prime}$ and edge sets $E^{\prime}$.
+Their natural extension to sets is $c(V^{\prime})=\sum_{v\in V^{\prime}}c(v)$ and $\omega(E^{\prime})=\sum_{e\in E^{\prime}}\omega(e)$ for vertex sets $V^{\prime}\subseteq V$ and edge sets $E^{\prime}\subseteq E$.
 The graph partitioning problem (GPP) asks for a partition of the vertex set $V$ into $k$ equally sized sets.
-More formally let $k\in\mathbb{N}$ be the desired number of blocks and let $\epsilon\geq 0$ be a hyper-parameter called imbalance.
-GPP then asks for a partition $V=V_{1}\cup\ldots\cup V_{k}$ such that $\forall i\neq j:V_{i}\cap V_{j}=\emptyset$ and each $V_{i}$ adheres to the balance constraint, that is $c(V_{i})\leq L_{\max}\coloneqq(1+\epsilon)\frac{c(V)}{k}$.
+More formally, let $k\in\mathbb{N}$ be the desired number of blocks and let $\epsilon\geq 0$ be a hyper-parameter called imbalance.
+GPP then asks for a partition $V=V_{1}\cup\ldots\cup V_{k}$ such that $\forall i\neq j:V_{i}\cap V_{j}=\emptyset$ and each $V_{i}$ adheres to the balance constraint, that is, $c(V_{i})\leq L_{\max}\coloneqq\lceil(1+\epsilon)\frac{c(V)}{k}\rceil$.
 Such a partition is called a $k$-way partition of $G$.
+We call a block $V_{i}$ overloaded if $c(V_{i})>L_{\max}$ and underloaded if $c(V_{i})\leq L_{\max}$.
 Let $E_{ij}=\{\{u,v\}\in E\mid u\in V_{i},v\in V_{j}\}$ be the edges crossing between block $V_{i}$ and $V_{j}$.
@@
 An entry $\mathcal{C}_{ij}=0$ denotes that the tasks do not communicate with each other.
-Most modern parallel applications only have a very sparse communication pattern between the tasks, which is why we model the communication matrix $\mathcal{C}$ as a task graph $G_{\mathcal{C}}$.
-For each non-zero entry $\mathcal{C}_{ij}$ the task graph has one forward and backward edge with weight $\mathcal{C}_{ij}$.
+Most modern parallel applications only have a very sparse communication pattern, which is why the application can also be modeled as a task graph $G_{\mathcal{C}}$.
+For each task in the application there exists one vertex with the same weight.
+For each non-zero entry $\mathcal{C}_{ij}$, the task graph has one forward and backward edge with weight $\mathcal{C}_{ij}$.
+It is assumed that the communication is symmetric, otherwise it can be modeled symmetrically [8].
 Let $k$ be the number of PEs on a supercomputer.
-The cost of communication between the PEs is captured by the hardware topology matrix $\mathcal{D}\in\mathbb{R}^{k\times k}$.
-An entry $\mathcal{D}_{xy}$ denotes a multiplicative cost factor that is applied to communication between PEs $x$ and $y$.
-The cost of task $i$ and $j$ communicating, if placed on PEs $x$ and $y$ respectively, is $\mathcal{C}_{ij}\mathcal{D}_{xy}$.
-The focus of this work is to tackle the general process mapping problem.
-It asks to assign the vertices of the task graph $G_{\mathcal{C}}$ to PEs of the supercomputer such that the amount of work on each PE is balanced and the total cost of communication is minimized.
-More formally, we ask for a mapping $\Pi:[n]\mapsto[k]$ such that $J(C,D,\Pi)\coloneqq\sum_{i,j}\mathcal{C}_{ij}\mathcal{D}_{\Pi(i)\Pi(j)}$ is minimized, while adhering to the balance constraint.
-In the hierarchical process mapping problem (HPMP), a special variant of the GPMP, the hardware topology of the supercomputer is described by a homogeneous hierarchy $H=a_{1}:\ldots:a_{\ell}$.
+The cost of communication between two PEs is captured by the hardware topology matrix $\mathcal{D}\in\mathbb{R}^{k\times k}$.
+An entry $\mathcal{D}_{xy}$ denotes a cost factor that is applied to communication between PEs $x$ and $y$.
+The cost of tasks $i$ and $j$ communicating, if placed on PEs $x$ and $y$ respectively, is therefore $\mathcal{C}_{ij}\mathcal{D}_{xy}$.
+It is assumed that all PEs can communicate with each other and that the matrix is symmetric.
+The general process mapping problem (GPMP) is then defined as follows.
+Let $G_{\mathcal{C}}=(V,E)$ with $n=|V|$ be a task graph, let $\mathcal{D}$ be the hardware topology matrix of a supercomputer with $k$ PEs and let $\epsilon$ be the allowed imbalance.
+GPMP then asks to partition the task set $V$ into $k$ equally sized subsets, such that the overall communication cost is minimized.
+More formally, determine a mapping $\Pi:[n]\mapsto[k]$ such that $J(\mathcal{C},\mathcal{D},\Pi)\coloneqq\sum_{i,j}\mathcal{C}_{ij}\mathcal{D}_{\Pi(i)\Pi(j)}$ is minimized, while each block adheres to the balance constraint.
+In the hierarchical process mapping problem (HPMP), a special variant of the GPMP, the hardware topology of the supercomputer is described by a hierarchy $H=a_{1}:\ldots:a_{\ell}$.
 It denotes that each processor has $a_{1}$ PEs, each node has $a_{2}$ processors, each rack has $a_{3}$ nodes, and so forth.
-In total, the system has $k=\prod_{1\leq\ell}a_{i}$ PEs.
+In total, the system has $k=\prod_{i\leq\ell}a_{i}$ PEs.
 Corresponding to the hierarchy, a distance $D=d_{1}:\ldots:d_{\ell}$ is given which denotes the cost factor between the PEs.
-Two PEs on the same processor have a cost factor of $d_{1}$, two PEs on the same node but on different processors have a cost factor of $d_{2}$ and so forth.
-In GPMP we typically assume $n>k$, i.e., we have (considerably) more tasks than PEs.
-If $n=k$ the problem is known as the one-to-one process mapping problem (OPMP) which is identical to the quadratic assignment problem (QAP).
-The QAP is commonly solved using refactorization and linearization techniques [3, 2, 21], of its corresponding mixed integer linear programming formulation [30].
+Two PEs on the same processor have a cost factor of $d_{1}$, two PEs on the same node but on different processors have a cost factor of $d_{2}$, two PEs on the same rack but on different islands have a cost factor of $d_{3}$, and so forth.
+The focus of this work lies in solving the HPMP.
+In GPMP, $n>k$ is typically assumed, i.e., there are (considerably) more tasks than PEs.
+If $n=k$, the problem is known as the one-to-one process mapping problem (OPMP) which is identical to the quadratic assignment problem (QAP).
+The QAP is commonly solved using refactorization and linearization techniques [3, 2, 22] of its corresponding mixed integer linear programming formulation [31].
 We refer the reader to [10, 5, 33] for comprehensive surveys.
-The problems GPP, GPMP, HPMP and QAP are all NP-hard problems [17, 39] and no constant factor approximation guarantee exists unless P = NP.
-II-B Multilevel Graph Partitioning
-Modern graph partitioning relies on the multilevel approach, which works in three phases, coarsening, partitioning and uncoarsening with refinement.
-In the coarsening phase a succession of smaller graphs $G,G_{1},\ldots,G_{t}$ is computed, with the goal that a smaller graph $G_{i+1}$ retains the same overall structure as $G_{i}$.
+The problems GPP, GPMP, HPMP, and QAP are all NP-hard problems [18, 40] and no constant factor approximation guarantee exists unless P = NP.
+2.1 Multilevel Graph Partitioning
+Modern graph partitioning relies on the multilevel approach, which works in three phases: coarsening, partitioning, and uncoarsening with refinement.
+In the coarsening phase, a succession of smaller graphs $G,G_{1},\ldots,G_{t}$ is computed, with the goal that a smaller graph $G_{i+1}$ retains the same overall structure as $G_{i}$.
 A “good” partition of $G_{i+1}$ would then also resemble a “good” partition for $G_{i}$.
-Coarsening is applied until the coarsest graph has less than $ck$ vertices, where $c$ is a tunable constant.
+Coarsening is applied until the coarsest graph has fewer than $ck$ vertices, where $c$ is a tunable constant.
+This small graph can then be efficiently partitioned.
+The contractions are then undone and the quality of the partition is improved.
 To coarsen a graph, one usually applies a matching or a clustering algorithm.
 The matching algorithm rates each edge and then determines a maximum weighted matching.
-Edge contraction is then applied on the determined matching.
-Edge contraction on an edge $\{u,v\}$ is applied by replacing vertices $u$ and $v$ with a new vertex $w$ with weight $c(w)=c(u)+c(v)$.
-Former neighbors of $u$ and $v$ are connected to $w$ with the same edge weight.
-Parallel edges are fused and their weight is summed.
+The edges selected in the matching are then contracted.
+Contracting an edge $\{u,v\}$ replaces the vertices $u$ and $v$ with a new vertex $w$ whose weight is defined as $c(w)=c(u)+c(v)$.
+All former neighbors of $u$ and $v$ are subsequently connected to $w$.
+Parallel edges are fused and their weights are summed.
 A clustering algorithm first determines clusters in the graph, and then contracts each cluster into one super-vertex.
@@
 In the initial partitioning phase, the coarsest graph $G_{t}$ is partitioned.
-Since $G_{t}$ is very small, expensive but high-quality algorithms, such as recursive bisection or direct $k$-way partitioning, can be applied to minimize edge-cut.
+Since $G_{t}$ is very small, expensive but high-quality algorithms, such as recursive bisection or direct $k$-way partitioning, can be applied to minimize the edge-cut.
 The uncoarsening and refinement phase iteratively uncontracts the graphs $G_{i+1}$ into $G_{i}$ by undoing the contractions.
@@
 The algorithms try to move vertices across the blocks to minimize the edge-cut, while respecting the balance constraint.
-Widely used refinement algorithms are the Fiduccia-Mattheyses (FM) algorithm [16], label propagation [38] and flow-based refinement [40].
+Widely used refinement algorithms are the Fiduccia-Mattheyses (FM) algorithm [17], label propagation [39], and flow-based refinement [41].
 Notable serial graph partitioners are Metis [28], KaFFPa [41], Jostle [49], Scotch [37] and KaHyPar [42].
-III Related Work
+3 Related Work
 Process mapping is closely related to graph partitioning, which has already seen a tremendous amount of research over the last few decades.
-We already gave a brief overview in Section II-B and refer the reader to [6, 10, 8, 42] for more comprehensive surveys.
-We will here give an overview of Jet [18], a GPU-based graph partitioning (Section III-A), and of process mapping (Section III-B).
-III-A GPU Multilevel Graph Partitioning
+We already gave a brief overview in Section 2.1 and refer the reader to [7, 11, 9, 43] for more comprehensive surveys.
+We will give here an overview of Jet [19], a GPU-based graph partitioner, and of process mapping.
+3.1 GPU Multilevel Graph Partitioning
 Gilbert et al.
@@
 Their main contribution includes a novel GPU-parallel refinement algorithm based on label propagation [39].
-They implemented their algorithm in the performance-portable Kokkos framework [14], enabling execution on GPUs and multi-core CPUs with a single code base.
+They implemented their algorithm in the performance-portable Kokkos framework [15].
 Matching.
 Jet utilizes a two-hop matching [30], originally developed for Mt-Metis [29], a shared-memory graph partitioner.
-The algorithm first applies a standard heavy-edge matching and only if more than 25% of vertices remain unmatched a two hop-matching is applied.
-A heavy edge-matching determines for each vertex its heaviest edge and two neighbors are matched if they agree on their heaviest edge.
+The algorithm first applies a heavy-edge matching and only if more than 25% of vertices remain unmatched, a two-hop matching is applied.
+A heavy-edge matching determines for each vertex its heaviest edge and two neighbors are matched if they agree on their heaviest edge.
 The two-hop matching can be split into three categories: leaf, twin and relative vertices.
@@
 To accelerate the matching of relatives, special matchmaker vertices with small degree are used.
-The two-hop matching first matches leaf vertices that share a neighbor, then twin vertices and in the end relative vertices, if necessary.
+At first leaf vertices that share a neighbor are matched, then twin vertices and at the end, relative vertices, if necessary.
 Coarsening.
-To coarsen the graph, a fine-grained per vertex hashing algorithm is deployed.
-First the maximum neighborhood size of each vertex in the coarse graph is overestimated.
-Then a prefix-sum over the sizes is determined, assigning each vertex a subarray to store its neighbors.
-Each vertex hashes its new neighboring vertices and inserts them and their weight into their respective subarray.
-In the end, the edges of all vertices are extracted into the classic Compressed Sparse Row (CSR) format.
+To coarsen the graph, a fine-grained per-vertex hashing algorithm is deployed.
+First, the maximum neighborhood size of each coarse vertex is overestimated.
+Then a prefix-sum over the sizes is determined, assigning each vertex a subarray.
+Each vertex hashes its new neighboring vertices and inserts them and their weights into their respective subarray.
+In the end, the edges of all vertices are extracted from their hash array.
 Initial Partitioning.
@@
 Additionally, while each individual vertex move might have a positive impact, moving adjacent vertices in parallel can worsen the solution quality.
-Gilbert et.
-al [18] tackle these problems by splitting the refinement into two parts, first unconstrained label propagation and afterwards rebalancing.
-The unconstrained label propagation starts by determining for each vertex $v$ the gain $F(v)$ (decrease in edge-cut) of moving it in an adjacent block.
+Gilbert et al.
+[19] tackle these problems by splitting the refinement into two parts, first unconstrained label propagation and afterwards rebalancing.
+The unconstrained label propagation starts by determining for each vertex $v$ the gain $F(v)$ (decrease in edge-cut) of moving it into an adjacent block.
 All potential vertex moves are then passed through the first filter $F(v)\geq 0\lor-F(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$.
 Here, $\operatorname{conn}(v,b)$ is the sum of edge weights from vertex $v$ to block $b$.
-To pass the first filter a move needs to be non-negative (fist condition) or its increase in edge-cut must be comparatively small (second condition).
-The constant $c\in[0,1]$ controls the number of negative gain moves.
-These moves are allowed to escape local minima.
-All moves that pass the first filter are collected in a list $\mathcal{X}$, which is implicitly sorted decreasingly by gain.
-For each vertex $v\in\mathcal{X}$, the algorithm then recomputes the gain $F_{2}(v)$ under an approximate future partition state: if a neighbor $u$ appears earlier in the ordering, it is assumed to have already moved to its new block.
-In a second filter only the vertices $v\in\mathcal{X}$ with a non-negative gain $F_{2}(v)$ are moved to their respective destination.
+To pass the first filter a move needs to be non-negative (first condition) or its increase in edge-cut must be comparatively small (second condition).
+The constant $c\in[0,1]$ controls the number of negative gain moves, with larger values allowing more moves.
+These negative moves are allowed to escape local minima.
+All moves that pass the first filter are collected in a list $\mathcal{X}$, which is implicitly sorted in decreasing order by gain.
+For each vertex $v\in\mathcal{X}$, the algorithm then recomputes the gain $\mathbb{F}(v)$ under an approximate future partition state: if a neighbor $u$ appears earlier in the ordering, it is assumed to have already moved to its new block.
+In a second filter, only the vertices $v\in\mathcal{X}$ with a non-negative gain $\mathbb{F}(v)$ are moved to their respective destination.
 To prevent oscillation each vertex that was moved is locked in the next iteration of unconstrained label propagation.
@@
 Jet implements two similar strategies, weak and strong rebalancing.
-Weak rebalancing determines for each vertex in an oversized partition $a$ the loss of moving it to a valid destination block.
+Weak rebalancing determines for each vertex in an overloaded block $A$ the loss of moving it to a valid destination block.
 Valid destination blocks are blocks $B$ with $c(B)\leq\sigma\leq L_{\max}$.
-The upper bound $\sigma$ is used instead of $L_{\max}$ to create a small buffer zone, so the block is not overloaded after only one move.
-For each oversized block $a$ its vertices are approximately sorted increasingly by loss in a list $L_{a}$.
+The upper bound $\sigma$ is used instead of $L_{\max}$ to create a small buffer zone, so a block is not overloaded after only one move.
+For each overloaded block $A$ its vertices are approximately sorted increasingly by loss in a list $L_{A}$.
 Then a prefix $L_{A}^{*}\subseteq L_{A}$ is determined such that moving all vertices in $L_{A}^{*}$ will balance block $A$.
-This minimizes the loss to balance a specific block, however, the destination blocks might become oversized, so another iteration might be necessary.
-Strong rebalancing reverses the process.
-Instead of oversized blocks pushing their vertices out, underloaded blocks will pull vertices from oversized blocks.
-They will only pull so many vertices such they themselves will not become oversized.
-This often achieves a balanced partition in just one iteration, however the loss is typically greater as in weak rebalancing.
-In practice, Jet applies two iterations of weak rebalancing followed by one iteration of strong rebalancing if necessary.
+This minimizes the loss to balance a specific block, however, the destination blocks might become overloaded, so another iteration might be necessary.
+Strong rebalancing takes extra care by preventing destination block from becoming overloaded.
+If this would happen, the vertices are moved to another underloaded block.
+However, these blocks must not necessarily be connected to the vertex.
+This results in a much greater loss but the partition will most likely be balanced.
 In total, Jet combines its unconstrained label propagation and rebalancing in multiple iterations and keeps the best solution encountered.
-The algorithm offers comparable solution quality to the well established CPU-based graph partitioner Metis [27], while being up to 20 times faster.
+The algorithm offers comparable solution quality to the CPU-based graph partitioner Metis [28], while being up to 20 times faster.
 Other GPU graph partitioners are Goodarzi et al.
 [20], Sphynx [1] and G-kway [32].
-III-B Process Mapping
-Similar to graph partitioning, process mapping has also seen a tremendous amount of research, and we refer the reader to Hoefler et.
-al [25] and Pellegrini [37] for broader overviews.
-Among the first to implement process mapping, was Hatazaki [22] who used graph partitioning to map MPI processes onto hardware topologies.
-Träff [46] and Yu et.
-al  [49] implemented mapping algorithms for special hardware, namely for the NEC SX-series of parallel vector computers and the Blue Gene/L Supercomputer, respectively.
-From a theoretical perspective it seems straightforward that minimizing $J(C,D,\Pi)$ will lead to a more performant algorithm.
-However, the questions remains if this is also true in practice.
-Brandfass et.
-al [7] showed that optimizing the communication cost indeed leads to performance improvements in the context of Computational Fluid Dynamics.
-Manwade et.
-al [33] show in a more general context that process mapping algorithms improve the performance of MPI applications.
-There are two main methods to solve process mapping, the two-phase approach and integrated mapping.
+3.2 Process Mapping
+Similar to graph partitioning, process mapping has also seen a tremendous amount of research, and we refer the reader to Hoefler et al.
+[26] and Pellegrini [38] for broader overviews.
+Among the first to implement process mapping was Hatazaki [23] who used graph partitioning to map MPI processes onto hardware topologies.
+Träff [47] and Yu et al.
+[50] implemented mapping algorithms for special hardware, namely for the NEC SX-series of parallel vector computers and the Blue Gene/L Supercomputer, respectively.
+From a theoretical perspective, it appears straightforward that minimizing $J(\mathcal{C},\mathcal{D},\Pi)$ will lead to a more efficient algorithm.
+However, the question remains whether the theoretical improvement translates into practical performance gains.
+Brandfass et al.
+[8] show that optimizing the communication cost indeed leads to performance improvements for the simulation of Computational Fluid Dynamics.
+Manwade et al.
+[34] show in a more general context that process mapping algorithms improve the performance of MPI applications.
+There are two main methods to solve process mapping: the two-phase approach and integrated mapping.
 The two-phase approach first leverages standard graph partitioning to determine a balanced $k$-way partition with small edge-cut of the task graph $G_{\mathcal{C}}$.
+This results in a partition with an overall small communication volume between the blocks.
 The second phase then determines a 1-to-1 mapping of blocks to the PEs that minimizes the communication cost.
-Müller-Merbach [35] presented a greedy algorithm to obtain an initial mapping, that served as a basis for many future algorithms.
-After the mapping phase a potential third refinement phase can be used to determine individual vertex moves that decrease the communication cost.
+Müller-Merbach [36] presented a greedy algorithm to obtain an initial mapping that served as a basis for many future algorithms.
+After the mapping phase, a potential third refinement phase can be used to determine individual vertex moves that decrease the communication cost.
 Heider [24] introduced a simple swap algorithm that considers all possible $\mathcal{O}(n^{2})$ swaps.
-This method was iteratively improved by Brandfass et.
-al [7] by avoiding redundant swaps, and Schulz and Träff [43], by utilizing better data structures and restricting the search space.
-Notable algorithms using the two-phase approach are GlobalMultisection [47] and SharedMap [44].
+This method was iteratively improved by Brandfass et al.
+[8] by avoiding redundant swaps, and Schulz and Träff [44] by utilizing better data structures and restricting the search space.
+Notable algorithms are GlobalMultisection [48] and SharedMap [45].
 GlobalMultisection [48] employs a method called hierarchical multisection inspired by KaFFPa-Map [44].
@@
 $G_{\mathcal{M}}$ is then perfectly partitioned alongside the hierarchy $H=a_{1}:\ldots:a_{\ell}$, first into an $a_{\ell}$-way partition, then each of the blocks into an $a_{\ell-1}$-way partition and so forth.
-This is recursively repeated until $k$ blocks are obtained which represent the $k$ PEs of the system.
+This is repeated recursively until $k$ blocks are obtained.
+These represent the $k$ PEs of the system.
 The mapping of blocks to PEs, and therefore vertices to PEs, can be determined by following the recursive calls.
 GlobalMultisection improves the approach by not creating the communication model graph $G_{\mathbb{M}}$, but instead directly determining an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ and then recursing on each block.
-Since no perfect balanced partition is needed, GlobalMultisection has more freedom to move vertices, resulting in lower communication costs.
+Since no perfectly balanced partition is needed, GlobalMultisection has more freedom to move vertices, resulting in lower communication costs.
 SharedMap [45] further decreases the communication cost.
 It introduces an adaptive imbalance strategy that rescales the imbalance parameter $\epsilon$ for each individual partitioning.
-It guarantees that the final $k$-way partition is $\epsilon$ balanced, and often gives each individual partitioning more space to work with.
+It guarantees that the final $k$-way partition is $\epsilon$-balanced.
 This results in an overall better mapping with lower communication costs.
-SharedMap uses KaFFPa [40] as one of its underlying graph partitioners, and it offers state-of-the-art solution qualities for process mapping.
+SharedMap uses the serial graph partitioner KaFFPa [41] as one of its underlying graph partitioners, and currently offers state-of-the-art solution quality for HPMP.
 The second approach to solve the GPMP, integrated mapping, combines the partitioning and mapping into one algorithm.
-It leverages the multilevel scheme and the refinement algorithms minimize $J(C,D,\Pi)$.
-Two prominent algorithm are IntegratedMapping [15] and Mt-KaHyPar-Steiner [24].
-IntegratedMapping [15] uses matching based coarsening to create a smaller representation of the graph.
+It leverages the multilevel graph partitioning scheme and the refinement algorithms minimize $J(\mathcal{C},\mathcal{D},\Pi)$ instead of edge-cut.
+Two prominent algorithms are IntMap [16] and Mt-KaHyPar-Steiner [25].
+IntMap [16] uses matching-based coarsening to create a smaller representation of the graph.
 The Global Path Algorithm [35], a $1/2$-approximation algorithm, is used to determine a matching of vertices.
@@
 As initial partitioning, the hierarchical multisection approach is used.
-Refinement strategies include quotient graph refinement [40], $k$-way FM refinement [40], label propagation [38] and multi-try FM refinement [40].
+Refinement strategies include quotient graph refinement [41], $k$-way FM refinement [41], label propagation [39], and multi-try FM refinement [41].
 All refinement methods use the gain of a vertex $v$ when moving it from its current block $\Pi(v)$ to a new block $b$.
 The gain is defined as follows:
-$G_{b}(v)=\sum_{u\in N(v)}C_{v,u}\Big(D_{\Pi(v),\Pi(u)}-D_{b,\Pi(u)}\Big).$ (1)
-Multiple methods are presented to query the distance $D_{i,j}$ between two PEs $i$ and $j$.
-The simplest one saves a matrix with $\mathcal{O}(k^{2})$ space and access each distance in $O(1)$.
-The other methods, have a considerable lower space complexity, but come with more computational load.
-Mt-KaHyPar-Steiner [24] originally was developed to minimize wire lengths in VLSI designs.
-Logical units, hypergraph $\mathcal{H}$, are mapped to chip locations, edge-weighted graph $\mathcal{C}$, such that the total sum of used edge-weights, the implicit wire-length distance, is minimized.
-Replacing $\mathcal{H}$ with the task graph $G_{\mathcal{C}}$ and $\mathcal{C}$ with the hardware topology matrix $D$ leads to GPMP.
+$G_{b}(v)=\sum_{u\in N(v)}C_{v,u}\Big(\mathcal{D}_{\Pi(v),\Pi(u)}-\mathcal{D}_{b,\Pi(u)}\Big).$ (1)
+Multiple methods are presented to query the distance $\mathcal{D}_{i,j}$ between two PEs.
+The simplest one saves a matrix with $\mathcal{O}(k^{2})$ space and accesses each distance in $O(1)$.
+The other methods, have a considerably lower space complexity, but come with more computational load.
+Mt-KaHyPar-Steiner [25] originally minimizes wire lengths in VLSI designs.
+Logical units, a hypergraph $\mathcal{H}$, are mapped to chip locations, an edge-weighted graph $\mathcal{C}$, such that the total sum of used edge weights, the implicit wire-length distance, is minimized.
+Replacing $\mathcal{H}$ with the task graph $G_{\mathcal{C}}$ and $\mathcal{C}$ with the hardware topology matrix $\mathcal{D}$ leads to GPMP.
 Mt-KaHyPar-Steiner is integrated into the shared-memory parallel hypergraph partitioner Mt-KaHyPar [21].
 The algorithm follows the multilevel (hyper-)graph partitioning approach.
-First the hypergraph is clustered, and these clustered are coarsed into a smaller graph.
+First, the hypergraph is clustered, and these clusters are contracted into individual vertices, creating a smaller graph.
 An initial $k$-way partition of the hypergraph is determined and then greedily mapped onto the chip locations.
-As local search refinement algorithms, label propagation [38], FM refinement [16] and flow-based refinement [40] are used.
-III-C Kokkos
-We utilize the Kokkos [14] library to implement our algorithms.
-Kokkos is a C++ library providing performance portability across many-core architectures, including GPUs.
+As local search refinement algorithms, label propagation [39], FM refinement [17], and flow-based refinement [41] are used.
+3.3 Kokkos
+We utilize Kokkos [15], a C++ library providing performance portability across many-core architectures, including GPUs.
 It abstracts two main concepts: (1) parallel execution patterns and (2) multidimensional arrays.
@@
 (a) parallel_for executes $N$ independent work units in parallel.
-Each work unit operates on the same user defined function $f(i)$ with only the index $i\in[0,N)$ changing for each work unit.
+Each work unit operates on the same user-defined function $f(i)$ with only the index $i\in[0,N)$ changing for each work unit.
 (b) parallel_reduce performs a reduction $R=\bigoplus_{i=0}^{N-1}f(i)$ over values returned by $f(i)$ using an associative operator $\oplus$.
-(c) parallel_scan computes an exclusive prefix sum over a range of values i.e.
+(c) parallel_scan computes an exclusive prefix sum over a range of values, i.e.
 $y_{i}=\sum_{j=0}^{i-1}f(j)$.
-All work units share access to memory, and Kokkos provides atomic operations and deterministic reduction mechanisms to prevent race conditions and ensure correctness on many-core architectures.
-III-D CSR Graph Format
-A graph $G=(V,E)$ is typically stored in the Compressed Sparse Row (CSR) format (also referred to as Compressed Row Storage - CRS).
+3.4 CSR Graph Format
+A graph $G=(V,E)$ is typically stored in the Compressed Sparse Row (CSR) format.
 The representation consists of an offset array $O$ (also called row-pointer) of size $|V|+1$ and two arrays of size $|E|$, one for the edge vertices $E_{v}$ and one for the edge weights $E_{w}$.
-The edges incident to vertex $v$ are stored in the contiguous range $[O[v],O[v+1]])$ of $E_{v}$ and $E_{w}$.
-Note that the graph is undirected and an edge $\{u,v\}$ is both stored for $u$ and $v$.
-This layout naturally supports parallelization across vertices, with each vertex processing its adjacency list independently.
-IV GPU-Accelerated Process Mapping
+The edges incident to vertex $v$ and the corresponding edge weights are stored in the contiguous range $[O[v],O[v+1])$ of $E_{v}$ and $E_{w}$.
+Note that the graph is undirected and an edge $\{u,v\}$ is stored for both $u$ and $v$.
+This layout naturally supports parallelization across vertices.
+4 GPU-Accelerated Process Mapping
 The main contribution of this work is the development of two GPU-based approaches for solving the hierarchical process mapping problem.
-Given the close relationship between process mapping and graph partitioning, and the recent success of GPU-based partitioners in achieving faster runtimes, we aim to extend these advantages to process mapping.
-Our first algorithm follows the hierarchical multisection approach.
-It recursively partitions the task graph $G_{\mathcal{C}}$ along the systems hierarchy: first across islands, then racks, and so forth.
+Given the close relationship between process mapping and graph partitioning, and the recent success of GPU-based partitioners in achieving faster runtimes, we extend these advantages to process mapping.
+The first algorithm follows the hierarchical multisection approach.
+It recursively partitions the task graph $G_{\mathcal{C}}$ along the system’s hierarchy: first across the islands, then across the racks, and so forth.
 The final mapping of blocks, and thus vertices, to PEs is obtained by following the recursive calls.
-The algorithm is presented in Section IV-A.
-The second algorithm integrates process mapping directly into a multilevel scheme and is inspired by the GPU-based graph partitioner Jet [18] (Section III-A).
+The algorithm is presented in Section 4.1.
+The second algorithm integrates process mapping directly into the multilevel graph partitioning scheme and is inspired by the GPU-based graph partitioner Jet [19].
 Its refinement phase has been adapted to employ the new gain equation (Equation 1).
-Further details are given in Section IV-B.
-IV-A Hierarchical Multisection
+Section 4.2 presents the algorithm.
+Extended CSR Format.
+Before presenting the algorithms, we want to present our extended CSR format.
+As introduced in Section 3.4, the CSR representation naturally supports parallelization over vertices, allowing each thread to process a vertex’s adjacency list independently.
+However, some computations benefit from direct parallelization over edges.
+In standard CSR, this requires nested parallel loops, first over vertices and then their neighbors.
+This can lead to poor load balancing due to varying vertex degrees, resulting in idle threads on GPUs.
+To overcome this, we extend the CSR format with an additional array $E_{u}$ of size $|E|$ that explicitly stores the opposite endpoint of each edge.
+This edge-centric representation enables flat parallelism over all edges, improving load balance and GPU utilization.
+We refer to this edge-list representation as $\mathbb{E}$ throughout the rest of this work.
+4.1 Hierarchical Multisection
 The hierarchical multisection approach was first introduced by Kirchbach et al.
@@
 First, an $a_{\ell}$-way partition of $G_{\mathcal{C}}$ is determined utilizing a graph partitioning algorithm.
-The resulting $a_{\ell}$ blocks, represent the largest hardware components of the supercomputer (e.g., the islands).
+The resulting $a_{\ell}$ blocks represent the largest hardware components of the supercomputer (e.g., the islands).
 Since the graph partitioner minimizes the edge-cut between the blocks, it directly corresponds to minimizing the communication cost across the islands.
@@
 This recursive procedure is continued until $k$ blocks are obtained.
-The mapping of the $k$ blocks to the $k$ PEs can be extracted by following the recursive calls.
 Schulz and Woydt [45] improved this method by introducing an adaptive imbalance parameter.
-Using the original imbalance parameter $\epsilon$ for each partitioning can overall result in an unbalanced $k$-way partition, i.e., some PEs may be overloaded.
+Using the original imbalance parameter $\epsilon$ for each partitioning can result overall in an unbalanced $k$-way partition.
 To avoid this problem, an adaptive imbalance parameter $\epsilon^{\prime}$ is determined for each partitioning.
@@
 Experimental results in [45] demonstrate that the adaptive imbalance improves the overall mapping quality.
-In particular, the SharedMap [44] implementation currently achieves state-of-the-art solution quality for the HPMP.
-To adopt the hierarchical multisection approach to GPUs, we must (1) employ a GPU-based graph partitioner and (2) construct the resulting subgraphs directly on the GPU.
-For the partitioning step, we use Jet [18] which we already presented in Section III-A.
-Jet outperforms CPU-based graph partitioners in terms of speed by factor in the range from 2 to 20 times while delivering equal or better solution quality.
+In particular, SharedMap [45] currently achieves state-of-the-art quality for the HPMP.
+To adopt the hierarchical multisection approach to GPUs, we must (1) employ a GPU-based graph partitioner to obtain the $a_{i}$-way partitions and (2) construct the resulting subgraphs directly on the GPU.
+For the partitioning step, we use our own reimplementation of Jet [19], which we already presented in Section 3.1.
+Jet outperforms the CPU-based graph partitioner Metis [28] in terms of speed by a factor in the range of 2 to 20 times while delivering equal or better solution quality.
 Alternative GPU-based graph partitioners such as Sphynx [1] or G-kway [32] could also be used.
-As no comprehensive comparison between GPU partitioners exists, we adopt Jet for its thorough evaluation.
-However, any faster or higher-quality alternative would translate to our algorithm.
-For the second step, we need to build the subgraphs of a graph $G=(V,E)$ given a partition $\Pi:[n]\mapsto[k]$.
-To circumvent the costly down- and upload between the CPU and GPU, the entire process is executed on the GPU.
-Our algorithm works in $k$ iterations, where each iteration builds one subgraph $G_{i}$.
+However, no comprehensive comparison between GPU partitioners exists, so we adopt the implementation of Jet for its thorough evaluation.
+Any faster or higher-quality graph partitioner would also improve this algorithm.
+For the second step, the induced subgraphs of a graph $G=(V,E)$ given a partition $\Pi:[n]\mapsto[k]$ need to be built.
+To circumvent the costly download and upload between the CPU and GPU, the entire process is executed on the GPU.
+The algorithm works in $k$ iterations, where each iteration builds one subgraph $G^{\prime}$.
+Algorithm 1 shows the pseudocode of one loop iteration.
+We define
+$\mathbf{1}[P]=\begin{cases}1&\text{if }P\text{ is true}\\ 0&\text{otherwise}\end{cases}$
+as an indicator function, which returns $1$ if the predicate $P$ is true and 0 else.
 Each iteration consists of three phases.
-The first phase determines the number of vertices $n_{i}$, edges and the subgraphs weight using three parallel_reduce loops.
-A vertex $v$ belongs to the $i$th subgraph if $\Pi(v)=i$ and likewise an edge $\{u,v\}$ belongs to it, if $\Pi(u)=\Pi(v)=i$.
-The second phase determines a mapping between the vertices of $G_{i}$ to the vertices of $G_{\mathcal{C}}$.
-The vertices of $G_{i}$ need to be indexed from 0 to $n_{i}-1$, which is also incorporated into the mappings.
-The mappings are needed so we can propagate the assigned partition of vertices in $G_{i}$ back to the vertices of $G_{\mathcal{C}}$.
-This step is efficiently implemented via a parallel_scan over the vertex set.
-The prefix sum assigns each vertex a unique new vertex in the range [0,n1)[0,n-1).
-In the third phase, we construct the subgraph using a parallel_scan.
-The scan first computes the CSR offsets for each vertex and then writes the corresponding edges and weights.
-For each vertex $u$ only edges $\{u,v\}$ in the same subgraph are inserted, while the mapping functions translate between global and local vertex IDs.
-Subgraph creation usually takes less than 2% of the total runtime of the complete algorithm.
-Algorithm 1 presents the pseudocode of the hierarchical multisection method.
-The function takes as input a (sub)graph $G_{L}$, the hierarchy $H$, the current hierarchy level $i$, an identifier list $I$, and a local mapping $\Pi_{L}$ that maps the vertices of $G_{L}$ to the vertices of $G_{\mathcal{C}}$.
-The identifier list $I$ uniquely identifies the position of the subgraph within the system hierarchy.
-Lines 242-4 handle the exit condition: when $i=0$, the subgraph represents one of the final $k$ blocks and all its vertices are assigned to exactly one of the $k$ PEs.
-The function $\textsc{calcID}(H,I)$ determines a unique ID for each identifier $I$ such that neighboring components receive consecutive IDs.
-The mapping $\Pi_{L}$ is used in Line 5 to translate the local vertices back to their original vertices in $G_{\mathcal{C}}$ and for these vertices the mapping is set to the ID.
-If the subgraph $G_{L}$ resides on a higher level ($i>0$) then further partitioning is required (Lines 6126-12).
-First, The adaptive imbalance $\epsilon^{\prime}$ is determined and then Jet [18] is utilized to partition $G_{L}$ into $a_{i}$ subgraphs via the mapping $\Pi^{\prime}:V(G_{L})\mapsto[a_{i}]$.
-For each resulting subgraph $G^{\prime}_{L}$, we determine a local mapping $\Pi^{\prime}_{L}$ from its vertices to $G_{\mathcal{C}}$ (needed in Line 5).
-The identifier list $I$ is copied and extended by index $j$ to uniquely identifier the subgraph (needed in Line 3).
-The algorithm then recurses on each subgraph.
-Finally, the complete algorithm is invoked in Line 14 with the task graph $G_{\mathcal{C}}$, the hierarchy $H$, the highest hierarchy level $\ell$, an empty identifier list, and the identity mapping for $G_{\mathcal{C}}$.
+The first phase determines the number of vertices $n^{\prime}$, edges $m^{\prime}$ and the subgraph’s weight $w^{\prime}$ using three parallel_reduce loops.
+A vertex $v$ belongs to the subgraph with id $k^{\prime}$ if $\Pi(v)=k^{\prime}$ and likewise an edge $\{u,v\}$ belongs to it if $\Pi(u)=\Pi(v)=k^{\prime}$.
+Lines 1 – 3 show this phase.
+The second phase remaps the vertices $v\in V_{i}$ from $[n]$ to $n^{\prime}]$.
+This remapping is a technical requirement, since graph partitioners expect a continuous range of vertices.
+The mapping $M:[n]\mapsto[n^{\prime}]$ is also later needed to propagate the resulting partition to the original vertices.
+Lines 585-8 show the pseudocode.
+Note that in Kokkos, a single parallel_scan is sufficient to construct the mapping.
+In the third phase, the subgraph is constructed.
+Lines 10 to 26 show the process.
+At first the new degree of each vertex is determined in parallel (Lines 10–13).
+Based on the new degrees the row pointer array of the graph can be determined using a prefix sum (Line 12).
+This assigns each vertex in the subgraph a continuous range, to place its edges.
+The variables $E^{\prime}$ in Lines 15–17 are the corresponding arrays of the subgraph.
+At last the actual insertion of the edges into the subgraph is done in Lines 18–26.
+For each vertex $v$ that belongs to the subgraph, its neighborhood is iterated over.
+For each neighbor $u$ that also belongs to the subgraph, the edge is inserted into the arrays and the index $i$ is incremented.
+Note that this whole phase can also be implemented as one large parallel_scan.
+Subgraph creation usually takes less than 5% of the total runtime of the hierarchical multisection algorithm.
+Algorithm 1 Build Subgraph
+1:Graph $G$, Partition $\Pi$, Subgraph ID $k^{\prime}$
+[…]
+[…]
+[…]
+5:
+[…]
+7:for $v\in V$ do in parallel
+8:  $B[v]\leftarrow\textbf{1}[\Pi(v)=k^{\prime}]$
+[…]
+10:
+[…]
+12:for $v\in V$ do in parallel
+13:  if $\Pi(v)=k^{\prime}$ then
+14:   $D[M(v)]\leftarrow\sum_{u\in N(v)}\textbf{1}[\Pi(u)=k^{\prime}]$
+[…]
+[…]
+[…]
+[…]
+19:for $v\in V$ do in parallel
+20:  if $\Pi(v)=k^{\prime}$ then
+21:   $i\leftarrow R^{\prime}[M(v)]$
+22:   for $u\in N(v)$ do
+23:     if $\Pi(u)=k^{\prime}$ then
+24:      $E^{\prime}_{v}[i]\leftarrow M(u)$
+25:      $E^{\prime}_{w}[i]\leftarrow\omega(u,v)$
+26:      $E^{\prime}_{u}[i]\leftarrow M(v)$
+27:      $i\leftarrow i+1$
+28:return $G^{\prime}=(R^{\prime},E^{\prime}_{v},E^{\prime}_{w},E^{\prime}_{u}),\ M$
+The complete hierarchical multisection algorithm can be seen in Algorithm 2.
+Here, it is implemented as a recursive function but an iterative approach would also be possible.
+The function takes as input a (sub)graph $G^{\prime}$, the hierarchy $H$ and the current hierarchy level $i$.
+Lines 2 and 3 handle the partitioning of the graph.
+At first the adaptive imbalance $\epsilon^{\prime}$ is computed using Equation 2 and afterwards our reimplementation of Jet [19] is used to partition the graph on the GPU.
+It determines a partition $\Pi^{\prime}$ into $a_{i}$ blocks with the specified $\epsilon^{\prime}$ imbalance.
+If this was the last partitioning of this recursive call ($i=1$) then the determined partition $\Pi^{\prime}$ needs to be propagated to the actual partition $\Pi$ of $G$.
+Since this is only a technical process we omit its details and only outline the brief idea.
+For each created graph $G^{\prime}$ keep a mapping $M^{\prime}$ from $V^{\prime}$ to the original vertices $V$ in $G$.
+Then for each $v\in V^{\prime}$ set $\Pi(M^{\prime}(v))=\Pi^{\prime}(v)$ into the final partition.
+The mapping $M^{\prime}$ can be computed by utilizing the mapping $M$ from Algorithm 1.
+If the subgraph $G^{\prime}$ resides on a higher level ($i>1$), then the resulting subgraphs need to be constructed and further partitioned.
+In Line 7 the subgraphs are built using Algorithm 1.
+Each subgraph is then recursively passed to the function, but with i1i-1 it is indicated that the subgraphs reside one level further down in the hierarchy.
+Finally, the complete algorithm is invoked in Line 10 with the task graph $G_{\mathcal{C}}$, the hierarchy $H$ and the highest hierarchy level $\ell$.
 Algorithm 2 Recursive Hierarchical Multisection
-1:Graph $G$, Hierarchy $\smash{H=a_{1}:\ldots:a_{\ell}}$
+1:Graph $G$, Hierarchy $H=a_{1}:\ldots:a_{\ell}$, Imbalance $\epsilon$
+2:function HM($G^{\prime},H,i$)
+3:  $\epsilon^{\prime}\leftarrow\textsc{AdapImb}(\ldots)$ $\triangleright$ uses Equation 2
+4:  $\Pi^{\prime}\leftarrow\textsc{GPU-GP}(G^{\prime},a_{i},\epsilon^{\prime})$ $\triangleright$ GPU Graph Partitioner
+5:  if $i=1$ then
+6:   propagate $\Pi^{\prime}$ to final partition $\Pi$
+7:   return
+8:  for $G^{\prime\prime}\in\textsc{CreateSubgraphs}(G^{\prime},\Pi^{\prime})$ do $\triangleright$ uses Algorithm 1
+9:   $\textsc{HM}(G^{\prime\prime},H,i-1)$}
+10:end function
 […]
-3:  if $i=0$ then
-4:   $\textsc{ID}\leftarrow\textsc{calcID}(H,I)$
-5:   for $\smash{v\in V_{L}}$ do $\smash{\Pi(\Pi_{L}(v))\leftarrow\textsc{ID}}$
-6:  $\smash{\epsilon^{\prime}\leftarrow\textsc{AdapImb}(\ldots)}$ $\triangleright$ uses Eq. 2
-7:  $\smash{\Pi^{\prime}\leftarrow\textsc{Jet}(G_{L},a_{i},\epsilon^{\prime})}$
-8:  $j\leftarrow 0$
-9:  for $G^{\prime}_{L}\in\textsc{CreateSubgraphs}(G_{L},\Pi^{\prime})$ do
-10:   $\Pi^{\prime}_{L}\leftarrow\textsc{MapToG}(G^{\prime}_{L},\Pi_{L})$
-11:   $\smash{\textsc{HM}(G^{\prime}_{L},H,i-1,I+[j],\Pi^{\prime}_{L})}$
-12:   $\smash{j\leftarrow j+1}$
-13:end function
-[…]
-IV-B Integrated Mapping
-The second main contribution of this work is an integrated mapping algorithm that incorporates the mapping directly into the graph partitioning.
-We follow the multilevel approach, but instead of minimizing the edge-cut we optimize $J(C,D,\Pi)$.
-The multilevel approach, coarsens the graph to a small representation, determines an initial partition and afterwards uncoarsens the graph and refines the mapping.
+4.2 Integrated Mapping
+The second main contribution of this work is a GPU-based integrated mapping algorithm that incorporates the mapping directly into the graph partitioning.
+We follow the multilevel approach, but instead of minimizing the edge-cut, we minimize $J(\mathcal{C},\mathcal{D},\Pi)$.
+Remember that the multilevel approach first coarsens the graph into a sequence of progressively smaller graphs.
+The smallest graph is then partitioned and the coarsening is reverted.
+Refinement algorithms are used during uncoarsening to further optimize the objective.
 We will detail our methodology for each phase in the next few paragraphs.
 We closely follow the matching, coarsening, refinement and rebalancing methods used by the GPU-based graph partitioner Jet [19] and adapt them to process mapping.
-Edge-List Graph Format.
-Before presenting the multilevel process mapping scheme, we want to present our extended CSR format.
-As introduced in Section III-D, the CSR representation naturally supports parallelization over vertices, allowing each thread to process a vertex’s adjacency list independently.
-However, some computations benefit from direct parallelization over edges.
-In standard CSR, this requires nested parallel loops, first over vertices, then their neighbors, which leads to poor load balancing due to varying vertex degrees, resulting in idle threads on GPUs.
-To overcome this, we extend the CSR format with an additional array $E_{u}$ of size $|E|$ that explicitly stores the opposite endpoint of each edge.
-This edge-centric representation enables flat parallelism over all edges, improving load balance and GPU utilization.
-We refer to this edge-list representation as $\mathbb{E}$ throughout the rest of this work.
 Matching.
-We utilize the two-hop matching presented by LaSalle et.
-al [29] but make a few changes.
+To determine a matching, the two-hop matching presented by LaSalle et al.
+[30] is adapted with a few minor changes.
 It starts with a heavy-edge matching algorithm.
-Instead of using $\omega(\{u,v\})$ as the weight of the edge, we use the rating function $\exp^{*2}(\{u,v\})=\frac{\omega(\{u,v\})^{2}}{c(u)c(v)}$ presented by Holtgrewe et al.
-[26] and also used in IntegratedMapping [15].
-Their experiments showed that it lead to improved running times and solution quality in the context of graph partitioning.
-By looping over $\mathbb{E}$ in parallel, we determine for each vertex $v$ its preferred neighbor $p(v)=u$ as the one with the greatest rating.
-Adjacent vertices $u$ with $c(v)+c(u)>L_{\max}$ are excluded to prevent an imbalanced initial solution.
-Afterwards, we determine in parallel for each vertex $v$ if $p(p(v))=v$, i.e., vertex $v$ prefers $p(v)$ and $p(v)$ prefers $v$.
-These vertices are matched to each other.
-If the vertices have multiple neighbors with the same rating, then it is very unlikely that $v$ and $p(v)$ prefer each other, and we would end up with few matches.
-We therefore add a small amount of deterministic noise $\eta(\{u,v\})$ onto each rating.
-The noise breaks the equality and is small enough to not influence the order of two distinct ratings.
-The matching procedure can be repeated multiple rounds, excluding previously matched vertices.
-Afterwards, leaf, twin and relative matchings are also applied multiple times if necessary.
+Instead of using $\omega(\{u,v\})$ as the rating of the edge, the rating function $\exp^{*2}(\{u,v\})=\frac{\omega(\{u,v\})^{2}}{c(u)c(v)}$ presented by Holtgrewe et al.
+[27] is used.
+Their experiments show that it leads to improved running times and solution quality in the context of graph partitioning.
+In parallel, each thread processes one vertex $v$.
+All unmatched neighbors are scanned and the vertex $u$ with the highest rating is saved as the preferred neighbor $p(v)=u$.
+To prevent identical ratings, a small amount of deterministic noise $\eta(\{u,v\})$ is added to each rating.
+It breaks any equalities, but is small enough to not influence the result.
+Vertices $u$ with $c(v)+c(u)>L_{\max}$ are excluded to prevent an imbalanced initial solution.
+Afterwards, each vertex $v$ is again processed in parallel.
+If $p(p(v))=v$, i.e., vertex $v$ prefers $p(v)$ and $p(v)$ prefers $v$, then $v$ and $p(v)$ are matched to each other.
+The matching procedure is repeated for multiple rounds until no more matches are determined.
+If more than $75\%$ of vertices are matched then the matching stops.
+Else, leaf, twin and relative matchings are applied to increase the number of matched vertices.
 We adopt the same hashing and matchmaker vertex strategy of Jet to improve performance.
-Instead of matching at least 75% of vertices, we reduce the percentage to 40%.
-While this leads to more levels in the multilevel framework, it also allows for more chances to refine the partition in the refinement phase.
-In our preliminary experiments using greater values lead to drastically worse solution quality.
-Coarsening.
-Edge-contraction, to create the coarsened graph, is based on a fine-grained per-vertex hashing scheme from Jet [18].
+In process mapping, matching-based coarsening approaches are usually advantageous, since the task graphs stem from scientific simulations and have a mesh-like structure.
+Another approach would be to cluster the graph, however this is usually superior in unstructured workloads like social networks.
+Contraction.
+Edge contraction, to create the coarse graph, is based on a fine-grained per-vertex hashing scheme from Jet [19], however, it is extended by using the extended CSR format.
+The overall idea is to assign each coarser vertex a hash array and the threads then hash the coarser edges into the respective array.
 The pseudocode can be seen in Algorithm 3.
-Let $G=(V,E)$ be the current graph and let $M:V\mapsto[n_{c}]$ map each vertex to its coarse counterpart.
-The value $n_{c}$ is the number of vertices in the coarse graph.
+Let $G=(V,E)$ be the current graph, $n_{c}$ the number of vertices in the coarse graph and $M:V\mapsto[n_{c}]$ map each vertex to its coarse counterpart.
 The algorithm first computes an upper bound on the neighborhood size of each coarse vertex by summing the degrees of all vertices assigned to it (Lines 131-3).
-In Line 4 and 5 two hash arrays, $H_{v}$ and $H_{w}$, are allocated that will hold the edges and their weight.
-The offset array $O$ denotes the index range $O[v]$ to $O[v+1]$ where the edges of the coarse vertex $v$ are placed.
-Next all edges $(u,v,w)\in\mathbb{E}$ are processed in parallel.
-If both endpoints contract to the same vertex, the edge becomes a self loop and is discarded (Line 8).
-Otherwise we need to insert $(M(v),w)$ into the adjacency of $M(u)$ in the interval $\mathcal{I}$.
+In Line 4 and 5 two hash arrays, $H_{v}$ and $H_{w}$, are allocated that will hold the edges and their weights.
+The offset array $O$ denotes the index range $[O[v],O[v+1])$ where the edges of the coarse vertex $v$ are placed.
+Then, all edges $(u,v,w)\in\mathbb{E}$ are processed in parallel.
+If both endpoints contract to the same vertex, the edge becomes a self-loop and is discarded (Line 8).
+Otherwise, we need to insert $(M(v),w)$ into the adjacency of $M(u)$ in the interval $\mathcal{I}$.
 Note that adding $(M(u),w)$ into the adjacency of $M(v)$ is handled by the work unit with reverse edge $(v,u,w)\in\mathbb{E}$.
 Insertion is done in Line 10, which we will explain in the next paragraph.
-After the parallel loop ends it remains to extract the actual graph from the arrays $H_{v}$ and $H_{w}$.
-This boils down to determining the true neighborhood sizes of each vertex and extracting the edges by skipping all Null entries in $H_{v}$.
-The most important part of our fine grained edge-contraction is the InsertOrLookup function in Line 10.
-To handle concurrent insertions without locks we utilize atomic compare-and-swap (CAS) operations.
-A CAS operation compares the value at $H_{v}[j]$ with Null and only if so swaps $M(v)$ in its place.
-Regardless if the check succeeded the previous value is returned.
-This means, if a thread executes a CAS operation on $H_{v}[j]$ and Null or $M(v)$ is returned, a slot for $M(v)$ was found.
-The weight $w$ will be atomically added to $H_{w}[j]$.
-Otherwise the thread will continue to cycle through the interval $\mathcal{I}$.
-Since we overestimated the neighborhood size of $M(u)$ at least one open slot will be available.
-To reduce atomic contention on $H_{v}[j]$, each thread determines $j\in\mathcal{I}$ by hashing $M(v)$ and then cycling through the range if necessary.
-Algorithm 2 GPU-based Coarsening
+After the parallel loop ends, it remains to extract the actual graph from the arrays $H_{v}$ and $H_{w}$.
+This boils down to determining the true neighborhood sizes for all vertices, building the offset and extracting the edges by skipping all Null entries in $H_{v}$.
+This can be accomplished by two parallel_scans.
+Algorithm 3 GPU-based Contraction
 1:Graph $G$, Mapping $M$, # Coarse Vertices $n_{c}$
@@
 3:for $v\in V$ do in parallel
-4:  $\smash{B[M(v)]\leftarrow B[M(v)]+|N(v)|}$
+4:  $B[M(v)]\leftarrow B[M(v)]+|N(v)|$}
 […]
@@
 9:  if $M(u)=M(v)$ then skip $\triangleright$ Self-loop discarded
-10:  $\smash{\mathcal{I}\leftarrow[O[M(u)],O[M(u)+1])}$
-11:  $\smash{\textsc{InsertOrLookup}(H_{v},H_{w},\mathcal{I},M(v))}$
-12:return $\smash{\textsc{ExtractCSR}(B,H_{v},H_{w})}$
+10:  $\mathcal{I}\leftarrow[O[M(u)],O[M(u)+1])$}
+11:  $\textsc{InsertOrLookup}(H_{v},H_{w},\mathcal{I},M(v))$}
+12:return $\textsc{ExtractCSR}(B,H_{v},H_{w})$}
+The most important part of the contraction is the InsertOrLookup function in Line 10.
+Each thread must place its coarse edge $(M(v),w)$ into $H_{v}$ and $H_{w}$ in the interval $\mathcal{I}$.
+In a serial algorithm one would first check if the edge $(M(v),w)$ already has an entry in $H_{v}$ and if so, add the weight to it.
+If there is no such entry then one would place the edge at a Null entry.
+However, in a parallel setting multiple threads add edges simultaneously and it must be ensured that edges are not overwritten and that duplicate edges are summed properly.
+To handle concurrent insertions, atomic compare-and-swap (CAS) operations can be utilized.
+A CAS operation $\text{CAS}(\textsc{ptr},\textsc{expected},\textsc{desired})$ compares the value at memory address ptr, with the expected vale and only if equal, writes the desired value in its place.
+Regardless of whether the thread wrote desired or not, the previous value stored at ptr is always returned.
+The important part of the CAS operation is that the check and the write are executed as one atomic operation.
+Two threads can never read the same expected value and write different desired values.
+A thread now inserts its edge by executing a CAS operation on memory address $H_{v}[j]$ with the expected value of Null and the desired value $M(v)$.
+If $M(v)$ is returned, then the edge was already written in $H_{v}[j]$ and the weight $w$ is atomically added to $H_{w}[j]$.
+If Null is returned, then this thread just wrote $M(v)$ into $H_{v}[j]$ and $w$ is also atomically added to $H_{w}[j]$.
+If neither is returned, then $H_{v}[j]$ already holds a different edge.
+The thread will continue cycling through the interval $\mathcal{I}$ until a valid position has been found.
+Since the neighborhood size of $M(u)$ was overestimated, at least one open slot will be available and the thread is able to place its edge.
+To reduce atomic contention on $H_{v}[j]$, each thread determines $j\in\mathcal{I}$ by hashing $M(v)$ and then cycling through the interval if necessary.
 Initial Partitioning.
-Initial partitioning is performed once the last coarsened graph has fewer than $128k$ vertices.
-Choosing a smaller threshold often resulted in an imbalanced initial partitions from which the algorithm failed to recover.
-We apply the hierarchical multisection approach explained in Section IV-A to acquire an initial mapping, but execute it entirely on the CPU.
+Initial partitioning is performed once the last coarsened graph has fewer than $8k$ vertices.
+We apply the hierarchical multisection [48, 45] approach, explained in Section 4.1, to acquire an initial mapping, but execute it entirely on the CPU.
 At this stage, the graph is sufficiently small that GPU execution offers no runtime advantage.
-Moreover, even the GPU-based partitioner Jet [18] would internally quickly fall back to the CPU-based partitioner Metis [27] when presented with such small graphs.
-To execute hierarchical multisection, we utilize the CPU-based partitioner KaFFPa [40] in its Fast configuration.
-Although the Strong configuration yields higher-quality mappings, its significantly increased runtime would dominate the total execution time, rendering the overall algorithm non-competitive.
+Moreover, even a GPU-based graph partitioner like Jet would internally quickly fall back to a CPU-based partitioner when presented with such small graphs.
+Also currently no GPU-based initial partitioning algorithms exist for HPMP or GP.
+To execute hierarchical multisection, we utilize a simple $k$-way graph partitioner.
+Uncontraction.
+Each coarser graph $G,G_{1},\cdots,G_{t}$ is separately stored in memory so uncontracting the graph from $G_{i+1}$ to $G_{i}$ is achieved by simply deleting $G_{i+1}$.
+The mapping $\Pi$, however, has to be updated.
+Let $M$ be the mapping that was used to create the coarser graph $G_{i+1}$ from $G_{i}$.
+Then set in parallel $\Pi(v)=\Pi(M(v))$ for each $v\in G_{i}$.
 Refinement.
-We adapt the unconstrained label propagation of Jet [18] to use $J(C,D,\Pi)$ as the objective function.
-The corresponding pseudocode is provided in Algorithm 3.
-The refinement algorithm begins by determining the best target block $\Pi^{\prime}(v)$ for each vertex $v$ by maximizing the gain $G_{b}(v)$ (see Equation 1) across all neighboring blocks (Lines 353-5).
+We adapt the unconstrained label propagation of Jet [19] to use $J(\mathcal{C},\mathcal{D},\Pi)$ as the objective function instead of edge-cut.
+Classic label propagation visits all vertices in a random order and if moving the current vertex to a neighboring block improves the objective and adheres to the balance constraint, then the move is made immediately.
+Therefore, the objective value never decreases and the balance constraint is always fulfilled.
+In a parallel setting, one would like to determine for each vertex in parallel whether moving it is beneficial.
+This part can easily be parallelized.
+Then, the beneficial moves would ideally also be executed in parallel.
+However, when moving the vertices in parallel, difficulties arise.
+Each individual move might adhere to the balance constraint, but moving multiple vertices at once can quickly overload a block.
+Also, although each individual move might be beneficial, moving adjacent vertices in parallel can have an overall negative effect on the solution quality.
+This makes it inherently difficult to parallelize label propagation.
+Jet’s approach works in two phases.
+At first an unconstrained label propagation algorithm is used that determines beneficial vertex moves.
+In this phase the balance constraint is completely ignored and subsequently a second rebalancing phase is required.
+This phase aims to restore the balance, while keeping the loss in the objective function small.
+The unconstrained label propagation works in two phases.
+At first, for each vertex the best move to a neighboring block is determined.
+These moves are filtered and only non-negative moves are considered further.
+Each move is then reevaluated under a future approximate partition state.
+If the move stays positive, then it is finally saved to be executed.
+The pseudocode of our adapted refinement algorithm is provided in Algorithm 4.
+The refinement algorithm begins by determining the best target block $\Pi^{\prime}(v)$ for each vertex $v$ by maximizing the gain $G_{b}(v)$ (see Equation 1) across all neighboring blocks (Lines 4 – 6).
+If indeed a move with non-negative gain has been found, then we insert the vertex into the list $\mathcal{X}$ via an atomically incremented index (Line 7).
 Algorithm 4 Unconstrained Label Propagation
-1:Graph $\smash{G}$, Mapping $\Pi$, Is-Locked array $L$
+1:Graph G{G}, Mapping $\Pi$
 […]
 […]
-4:for $v\in V$ do in parallel
-5:  if $\smash{L[v]=1}$ then skip $\triangleright$ Vertex locked
-6:  $\smash{B\leftarrow\{\Pi(u)\mid u\in N(v)\}}$
-7:  $\smash{\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)}$
-8:  if $\smash{G_{\Pi^{\prime}(v)}(v)\geq 0}$ then $\smash{\mathcal{X}[v]=1}$
 […]
-10:for $v\in V$ do in parallel
-11:  if $\smash{\mathcal{X}[v]=1\text{ and }\smash{\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0}}$ then $\smash{\mathcal{M}[v]=1}$
+5:for $v\in V$ do in parallel
+6:  $B\leftarrow\{\Pi(u)\mid u\in N(v)\}$}
+7:  $\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)$}
+8:  if $G_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{X}[i++]=v$}
 […]
-13:return $\smash{\mathcal{M},\Pi^{\prime},L}$
-In Jet, a filter is applied that retains only vertex moves satisfying $G_{b}(v)\geq 0\lor-G_{b}(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$.
-This criterion is suitable when minimizing the edge-cut, as both $G_{b}(v)$ and $\operatorname{conn}(v,\Pi(v))$ depend solely on edge weights and are thus directly comparable.
-In contrast, for process mapping, $G_{b}(v)$ contains the additional communication factors and is therefore often much smaller than $\operatorname{conn}(v,\Pi(v))$ and varies significantly, depending on $v$’s connections.
-This renders the original condition ineffective.
-We therefore only allow non-negative moves $G_{b}(v)\geq 0$ to pass the filter, as it performed best in our preliminary experiments.
+[…]
+11:for $v\in\mathcal{X}$ do in parallel
+12:  if $\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{M}[i++]=v$}
+13:return $\mathcal{M},\Pi^{\prime}$}
+Originally, Jet applies a more relaxed filter that also retains negative vertex moves.
+A vertex $v$ must only satisfy $G_{b}(v)\geq 0\lor-G_{b}(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$.
+The value $\operatorname{conn}(v,\Pi(v))$ denotes the sum of edge weights connecting $v$ and block $\Pi(v)$.
+Negative vertex moves are allowed to pass this filter, such that the algorithm may escape local minima.
+The criterion is suitable when minimizing the edge-cut, as both $G_{b}(v)$ and $\operatorname{conn}(v,\Pi(v))$ depend solely on edge weights and are thus directly comparable.
+In contrast, for process mapping, $G_{b}(v)$ contains the additional communication factors and is therefore often much greater than $\operatorname{conn}(v,\Pi(v))$.
+This renders the original filter ineffective.
+Therefore only non-negative moves $G_{b}(v)\geq 0$ are allowed to pass the filter and be inserted into $\mathcal{X}$, as it performed best in our preliminary experiments.
 A more sophisticated filtering criterion may further improve solution quality and remains a topic for future work.
-In Line 7, vertices passing this filter are inserted into the set $\mathcal{X}$, which is implemented as a binary array marking selected vertices.
-The set is implicitly ordered by $\operatorname{ord}$, defined as
-$\operatorname{ord}(u)\begin{cases}<\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)>G(v),\\ <\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G(u)=G(v)\land u<v,\\ >\operatorname{ord}(v)&\text{otherwise}.\end{cases}$
-Next, all vertex moves in $\mathcal{X}$ are reevaluated under an approximated future mapping state.
-If a vertex $v$ has a neighbor $u$ scheduled to move to block $\Pi^{\prime}(u)$ with $\operatorname{ord}(u)<\operatorname{ord}(v)$, then $u$ is assumed to have already moved when recomputing the gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$.
-We mark each vertex $v\in\mathcal{X}$ in a To-Move array $\mathcal{M}$ if it has non-negative gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ (Lines 9109-10).
-To prevent oscillation of vertices, the To-Move array serves as the locked array $L$ for the next iteration (Line $11$).
-The function returns the To-Move array $\mathcal{M}$ the destination mapping $\Pi^{\prime}$ and the Locked array $L$.
+The list $\mathcal{X}$ now holds all non-negative vertex moves.
+In a serial algorithm, one could now sort the list and execute the best vertex moves first.
+Moves further to the end of the list could then be reevaluated under the new partition state and only be executed if they are positive.
+However, in a parallel setting this is not straightforward.
+Firstly, sorting on a GPU is expensive and should be avoided and secondly, executing one move after the other is inherently serial.
+Gilbert et al.
+[19] therefore first only implicitly sort the list $\mathcal{X}$ and then each vertex move is re-evaluated under an approximate future partition state in parallel.
+The list $\mathcal{X}$ is unordered, but we can define an ordering on its elements via
+$\operatorname{ord}(u)\begin{cases}<\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G_{\Pi^{\prime}(u)}(u)>G_{\Pi^{\prime}(v)}(v),\\ <\operatorname{ord}(v)&\text{if }u\in\mathcal{X}\land G_{\Pi^{\prime}(u)}(u)=G_{\Pi^{\prime}(v)}(v)\land u<v,\\ >\operatorname{ord}(v)&\text{otherwise}.\end{cases}$
+Vertices with a high gain are at the front of the list, while vertices with lower gain are towards the end of the list.
+By querying $G_{\Pi^{\prime}(v)}(v)\geq 0$} we can determine whether a vertex was inserted into $\mathcal{X}$.
+It is now assumed that vertices with a lower ordering would have moved in a serial algorithm before a vertex with a higher ordering.
+All vertex moves in $\mathcal{X}$ are now reevaluated in parallel under an approximate future mapping state.
+We denote the reevaluated gain by $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ and it is determined just like the gain $G_{\Pi^{\prime}(v)}(v)$ (Equation 1), only the assigned block of some neighbors change.
+If a vertex $v$ has a neighbor $u$ with $\operatorname{ord}(u)<\operatorname{ord}(v)$ and vertex $u$ is scheduled to move to block $\Pi^{\prime}(u)$, then $u$ is assumed to have already moved when recomputing the gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$.
+Each vertex $v\in\mathcal{X}$ that has a non-negative gain $\mathbb{G}_{\Pi^{\prime}(v)}(v)$ is then added to the final list of moves (Lines 9 – 11).
+To prevent oscillation of vertices, all vertices in the final list $\mathcal{M}$ are locked for the next iteration of label propagation.
+The results are also cached and if the neighborhood of a vertex did not change, then its result is reused.
+Both of these techniques are not shown in Algorithm 4 for visual clarity.
+The function returns the list $\mathcal{M}$ and the destination mapping $\Pi^{\prime}$.
 Rebalancing.
-Our rebalancing is also adapted from Jet [18].
-Jet splits its rebalancing into a weak and strong algorithm.
-The weak rebalancing keeps the loss small, however, it might need multiple iterations to achieve a balanced partition.
-The strong rebalancing instead often achieves a balanced partition in one iteration, but has a greater loss.
-Weak rebalancing starts by determining for each vertex $v$ in an overloaded block, the move with maximum gain (or equally minimum loss) $G_{b}(v)$ into a neighboring block $b$ that has a weight smaller than the threshold $\sigma\leq L_{\max}$.
-The threshold $\sigma$ is chosen such that there is a small dead-zone between $\sigma$ and $L_{\max}$, so $b$ can take multiple vertices without instantly being overloaded.
-In our implementation we start with $\sigma=L_{\max}-0.065L_{\max}$ and gradually lower it throughout the multilevel scheme until at the lowest level we use $\sigma=L_{\max}-0.005L_{\max}$.
-We lower it, since the vertex weights are also gradually getting smaller during uncoarsening.
-Therefore, we can operate on a more fine level and more blocks become eligible.
-If for a vertex no such neighboring block exists, a random block that fits the threshold is chosen.
-Ideally, each block would store all its outgoing vertex moves decreasingly in a list and then a prefix of the list could be chosen, such that the block becomes balanced.
-This would grant the smallest possible loss based on the current partition.
-However, since sorting can be expensive on a GPU, we utilize an approximation of the sorted list similar to Jet.
-Each vertex move is assigned a slot in the list $0,-1,\ldots,-10,-20,\ldots,-100,-200,\ldots,-1000,\operatorname{X}]$.
-The slot ++ holds all positive gain moves, the slot $\operatorname{X}$ contains all very large negative gain moves and all other slots hold moves with gain $L[i]\geq G_{b}(v)>L[i+1]$.
-Note that at first we increase the range by 1, then by 10 and in the end by 100.
-This gives us an increased resolution for moves with a comparatively small loss, which minimizes the loss when only a few vertices need to be moved.
-Note that Jet also adopts the ++ and 0 slot and for any negative moves they use $\log_{2}(-G_{b}(v))$ to assign a vertex a slot.
-The approximated list is build in four phases, first by counting how many vertices fall into each slot, allocating enough memory, building a exclusive prefix-sum over the sizes and in the last phase inserting the vertices into their slots.
-To reduce atomic contention we utilize $\rho=2$ mini buckets in each slot and a vertex $v$ is assigned its mini bucket by $v\mod\rho$.
-As the last step of weak rebalancing a list prefix is chosen for each block $b$ and all vertices in it are moved.
-The prefix moves at least $c(b)-L_{\max}$ weight, so that the block $b$ becomes balanced.
-Since the lists are approximately sorted by their gain, the prefixes have a small combined loss.
+The rebalancing is also adapted from Jet [19], which splits its rebalancing into a weak and strong part.
+The idea is to determine for each vertex in an overloaded block the move with minimal loss to an underloaded block.
+All moves are approximately sorted and then as few vertices as necessary are moved.
+Weak rebalancing aims to keep the loss small, but might need multiple iterations, while strong rebalancing balances in one iteration, but has a much greater loss.
+The weak rebalancing starts by determining for each vertex $v$ in an overloaded block the move with maximum gain (or equally minimum loss) $G_{b}(v)$ into a neighboring block $B$ that has a weight less than the threshold $\sigma\leq L_{\max}$.
+The threshold $\sigma$ is chosen such that there is a small dead zone between $\sigma$ and $L_{\max}$, so $B$ can take multiple vertices without instantly being overloaded.
+In our implementation we set $\sigma=L_{\max}-100$, just like Jet.
+If for a vertex no such neighboring block exists, a random block that satisfies the threshold is chosen.
+Heavy vertices $v$ that weigh more than $1.5(c(\Pi(v))-c(V)/k)$ are not allowed to move during rebalancing.
+Ideally, each block would store all its outgoing vertex moves in decreasing order in a list and then a prefix of the list could be chosen, such that the block becomes balanced.
+This would grant a relatively small loss based on the current partition.
+However, sorting on a GPU is expensive, so instead an approximation of the sorted list is utilized.
+Each negative vertex move is assigned a slot in a $\log_{2}$–spaced bucket list.
+A vertex $v$ with negative gain $G_{b}(v)$ would therefore be stored in bucket $i$ with $i\leq\log_{2}(-G_{b}(v))<i+1$.
+This gives an increased resolution for moves with a comparatively small loss, which minimizes the loss when only a small number of vertices need to be moved.
+Strictly positive vertex moves are assigned to a special ++-bucket and 0-gain moves are placed in a special 0-bucket.
+Instead of first building the approximation list explicitly and then determining the prefix of vertex moves, one can also use a per-vertex decision process.
+For each vertex, when it is added to a bucket it will store the accumulated weight of vertices already inside of its bucket.
+After all vertices have been processed, a prefix sum over all bucket weights is computed.
+Through this, each vertex can now determine the sum of the vertex weights in earlier buckets and it knows how much vertex weight will be moved in its own bucket, before the vertex itself is moved.
+If this weight is smaller than $c(B)-L_{\max}$, then this vertex needs to be moved.
+Algorithm 5 Weak Rebalancing
+1:Graph $G=(V,E)$, Mapping $\Pi$
+[…]
+3:for $v\in V$ do in parallel
+4:  if $c(\Pi(v))\leq L_{\max}$ or $v$ too heavy do Skip
+5:  $B\leftarrow\{\Pi(u)\mid u\in N(v)\land c(\Pi(u))\leq\sigma\}$}
+6:  if $B=\emptyset$ then
+7:   $B\leftarrow\{\text{rnd Block b with }c(b)\leq\sigma\}$
+8:  $g\leftarrow\max_{b\in B}G_{b}(v)$
+9:  $\Pi^{\prime}(v)\leftarrow\text{argmax}_{b\in B}G_{b}(v)$}
+[…]
+[…]
+12:for $v\in V$ do in parallel
+13:  if $\mathbb{G}_{\Pi^{\prime}(v)}(v)\geq 0$} then $\mathcal{M}[i++]=v$}
+14:return $\mathcal{M},\Pi^{\prime}$}
+By approximately sorting the lists by gain, and only moving as few vertices as possible, the overall combined loss stays small.
 However, there is no safeguard against the destination blocks becoming overloaded.
 While the threshold $\sigma$ helps in preventing this, it does not guarantee it.
-Therefore, the destination blocks might become overbalanced and another iteration is required.
-Strong rebalancing operates in a similar manner to weak rebalancing but reverses the assignment perspective.
-As before, each vertex $v$ in an overloaded block $b$ identifies an eligible neighboring target block $b^{\prime}$ that minimizes the loss of moving $v$.
-However, instead of storing the vertex moves with the source block $b$, they are stored in the lists of the corresponding target block $b^{\prime}$.
-Each target block then selects a prefix of moves from its own list, such that it stays balanced after accepting them.
-After execution, the target blocks are guaranteed to stay balanced, while the source blocks lose weight and are also likely to be balanced.
-Consequently, strong rebalancing can restore balance in just one iteration.
-However, target blocks will often move more vertices than necessary, resulting in a higher overall loss compared to weak rebalancing.
+Therefore, the destination blocks might become overloaded and another iteration is required.
+The strong rebalancing operates in a very similar manner to weak rebalancing but takes extra precaution to not overload the destination blocks.
+Vertices that would overload a destination block are reassigned to other underloaded blocks.
+However, the vertices are not necessarily connected to their new destination block, which results in a much greater loss.
+We adapted the method to use $J(\mathcal{C},\mathcal{D},\Pi)$ as the objective function.
+Although this increases computational complexity, we observed no improvement in solution quality.
+Therefore, we use the standard rebalancing that minimizes edge-cut, which performs equally well but is faster to compute.
+One explanation is that edge-cut loss and communication cost loss correlate, so moves with small edge-cut loss also incur small communication cost loss.
+So regardless of the objective function, similar sets of vertices are moved.
+Additionally, any poor rebalancing move is likely corrected during the label propagation phase due to its high reversal gain.
 Algorithm 6 Overall Refinement Algorithm
@@
 […]
-[…]
-5:while $i<i^{\max}$ do
+4:while $i<12$ do
+5:  $i\leftarrow i+1$
 6:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
-7:   $\mathcal{M},$ $\Pi^{\prime\prime}$, $L$ $\leftarrow\operatorname{UnconLabelProp}(G,\Pi^{\prime},L)$
+7:   $\mathcal{M},$ $\Pi^{\prime\prime}\leftarrow\operatorname{UnconLabelProp}(G,\Pi^{\prime})$
 8:   $i_{w}\leftarrow 0$
 9:  else
-10:   $L\leftarrow\textsc{Zero}(|V|)$ $\triangleright$ Reset Locks
-11:   if $i_{w}<i^{\max}_{w}$ then
-12:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{weakReb}(G,\Pi^{\prime})$
-13:     $i_{w}\leftarrow i_{w}+1$
-14:   else
-15:     $\mathcal{M},$ $\Pi^{\prime\prime}$ $\leftarrow\operatorname{strongReb}(G,\Pi^{\prime})$
-16:     $i_{w}\leftarrow 0$
-17:  $\Pi^{\prime}\leftarrow\operatorname{move}(\mathcal{M},\Pi^{\prime\prime})$
-18:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
-19:   if $J(C,D,\Pi^{\prime})<J(C,D,\Pi)$ then $\Pi\leftarrow\Pi^{\prime}$
-20:   if $J(C,D,\Pi^{\prime})<\phi J(C,D,\Pi)$ then $i\leftarrow 0$
-21:  else
-22:   if $\operatorname{maxImb}(G,\Pi^{\prime})<\operatorname{maxImb}(G,\Pi)$ then
-23:     $\Pi\leftarrow\Pi^{\prime},\quad i\leftarrow 0$
+10:   if $i_{w}<2$ then
+11:     $\mathcal{M},\Pi^{\prime\prime}\leftarrow\operatorname{weakReb}(G,\Pi^{\prime})$
+12:     $i_{w}\leftarrow i_{w}+1$
+13:   else
+14:     $\mathcal{M},\Pi^{\prime\prime}\leftarrow\operatorname{strongReb}(G,\Pi^{\prime})$
+15:     $i_{w}\leftarrow 0$
+16:  $\Pi^{\prime}\leftarrow\operatorname{Move}(\mathcal{M},\Pi^{\prime\prime})$
+17:  if $\operatorname{maxImb}(G,\Pi^{\prime})\leq L_{\max}$ then
+18:   if $J(\mathcal{C},\mathcal{D},\Pi^{\prime})<J(\mathcal{C},\mathcal{D},\Pi)$ then $\Pi\leftarrow\Pi^{\prime}$
+19:   if $J(\mathcal{C},\mathcal{D},\Pi^{\prime})<\phi J(\mathcal{C},\mathcal{D},\Pi)$ then $i\leftarrow 0$
+20:  else
+21:   if $\operatorname{maxImb}(G,\Pi^{\prime})<\operatorname{maxImb}(G,\Pi)$ then
+22:     $\Pi\leftarrow\Pi^{\prime},\quad i\leftarrow 0$
 Overall Refinement Algorithm.
 The complete refinement procedure is summarized in Algorithm 6.
-The algorithm iteratively alternates between unconstrained label propagation and rebalancing (depending on the current imbalance) until the maximum number of iterations $i^{\max}$ is reached.
-A copy of the mapping $\Pi^{\prime}$ is used and the actual mapping $\Pi$ is only updated if an improvement mapping has been found.
-The parameter $i^{\max}_{w}$ controls how many consecutive iterations of weak rebalancing are applied before a single iteration of strong rebalancing is triggered.
+The algorithm iteratively alternates between unconstrained label propagation and rebalancing (depending on the current imbalance) until at least 12 iterations have been executed.
+This guarantees that at least three full cycles of label propagation, then two weak rebalancing steps and at last one strong rebalancing can be executed.
+A copy $\Pi^{\prime}$ is used and the actual mapping $\Pi$ is only updated if an improved mapping has been found.
 In each iteration, if the current partition $\Pi^{\prime}$ satisfies the imbalance constraint (Line 5), the algorithm performs unconstrained label propagation to reduce the communication cost (Line 6).
-Otherwise, the algorithm invokes weak or strong rebalancing to improve the balance (Lines 8158-15).
+Otherwise, the algorithm invokes weak or strong rebalancing to improve the balance (Lines 8 – 15).
 After each iteration, the partition $\Pi^{\prime}$ is updated (Line 16), and the best feasible solution encountered so far is saved.
 The process terminates when no significant improvement in either objective quality or balance has been achieved within the iteration limits.
-The hyperparameter $\phi=0.999$ specified the minimum relative improvement of $J(C,D,\Pi)$ that has to be achieved in each iteration.
-We set $i^{\max}_{w}=2$ by default, but increase it to $10$ in the final refinement stage of the multilevel pipeline to ensure minimal loss while achieving balance.
-The maximum number of iterations $i^{\max}$ is set to $12+\operatorname{level}$, where $\operatorname{level}$ denotes the current level in the coarsening hierarchy (the original graph at level 0, the coarser representation at higher levels).
-At coarser levels, graphs are smaller, allowing more iterations, but at finer levels, graphs are larger, so each iteration is more computationally expensive.
-Additional Data Structure.
-Determining the gain of moving a vertex to its neighboring blocks is essential for both label propagation and rebalancing.
-While this information can be obtained by iterating over the neighbors of a vertex, it can be inefficient as a vertex with 100 neighbors might only be connected to two blocks.
-We employ an additional data structure [18] that enables rapid access to the neighboring blocks.
-Conceptually, it is similar to the CSR representation of a graph, however, instead of storing edges, it stores the neighboring blocks and their the aggregated weight of edges.
-For each vertex $v$, we initially allocate memory for $|\{\Pi(u)\mid u\in N(v)\}|+2$ entries.
-The additional capacity accounts for potential future vertex moves.
-Note that this is less memory than Jet, which allocates $\min\{k,|N(v)|\}$ per vertex.
-Whenever a vertex changes its block assignment, we attempt to update its connectivity within the existing memory.
-We update the connectivity of $u$ to its neighbors $v$ as follows.
-For each $(u,v,w)\in\mathbb{E}$, we subtract $w$ from $v$’s connection to $\Pi(u)$.
-Should the resulting weight drop to 0, then the slot is freed.
-Next, we add the connectivity to $u$’s new block $\Pi^{\prime}(u)$ in $v$’s connectivity, if $\Pi^{\prime}(u)$ does not already exist.
-If this is not possible, because all slots are already occupied, we mark in a bit array, that $\Pi^{\prime}(u)$ could not be inserted into $v$’s adjacency.
-As a last step we add the weights $w$ to the slot $\Pi^{\prime}(u)$.
-All these steps utilize a parallel_for loop over the edge list $\mathbb{E}$.
-If any vertex cannot be updated because not enough slots are available, we trigger a full rebuild of the data structure.
-Any vertex that exceeded its capacity receives $|\{\Pi(u)\mid u\in N(v)\}|+2$ slots, while all other vertices retain their slots.
-V Experimental Evaluation
-All of our algorithms111github.com/HenningWoydt/GPU-HeiProMap are implemented in C++ 17 and utilize the Kokkos library [14] version 4.7.0 for the GPU code.
-It compiles the host side code with GCC 11.4 and the GPU code with Nvidia Cuda compiler 12.6.
-We use KaFFPa [40] from the KaHIP library version 3.19.0, which is also implemented in C++.
+The hyperparameter $\phi=0.999$ specifies the minimum relative improvement of $J(\mathcal{C},\mathcal{D},\Pi)$ that has to be achieved in each iteration.
+The constants 12, 2 and $\phi$ are transferred from Jet.
+During the algorithm, an additional structure (not shown in the pseudocode) stores for each vertex $v$ all neighboring blocks and the sum of edge weights to those blocks.
+This structure is used to efficiently compute gains during label propagation and rebalancing.
+For each vertex it holds a hash array of size $\min(|N(v)|,k)$ to store all possible connected blocks.
+In contrast to Jet, the structure is initialized by using the extended CSR format.
+By using an edge-parallel loop each thread cycles through the vertex hash array and uses CAS operations to find a valid entry.
+When moving the vertices (Line 15) the structure must be updated.
+This is done by either completely resetting the affected vertices’ hash arrays and filling them from scratch, or by first removing old connections and then adding new connections.
+5 Experimental Evaluation
+All of our algorithms222 / GPU-HeiPa are implemented in C++ 20 and utilize the Kokkos library [15] version 5.0.0 for the GPU code.
+It compiles the host-side code with GCC 11.4.0 and the GPU code with NVIDIA Cuda compiler 12.6.77.
 We compile with the highest optimization (-O3).
 System.
-The experiments are conducted on a machine that has an Intel® Xeon® w5-3435X, that runs on a max frequency of 3.13.1Ghz.
+The experiments are conducted on a machine that has an Intel® Xeon® w5-3435X that runs at a max frequency of $3.1$ GHz.
 The system features 128 GB of main memory.
-The used GPU is a NVIDIA® GeForce RTX™ 4090 with 24GB of memory.
-It has 16.384 Cuda cores with a boost frequency of 2.52.5Ghz.
+The GPU is an NVIDIA® GeForce RTX™ 4090 with 24 GB of memory.
+It has 16,384 Cuda cores with a boost frequency of $2.5$ GHz.
 The system runs Ubuntu 22.04.4.
+Table 1: Benchmark instances with their corresponding numbers of vertices and edges.
+SuiteSparse Other Walshaw
+Graph $|V|$ $|E|$ Graph $|V|$ $|E|$ Graph $|V|$ $|E|$
+cop20k_A 99 843 $\approx$1.3M afshell9 504 855 $\approx$8.5M 598a 110 971 741 934
+2cubes_sphere 101 492 772 886 thermal2 $\approx$1.2M $\approx$3.7M fe_ocean 143 437 409 593
+thermomech_TC 102 158 304 700 nlr $\approx$4.2M $\approx$12.5M 144 144 649 $\approx$1.1M
+cfd2 123 440 $\approx$1.5M deu $\approx$4.4M $\approx$5.5M wave 156 317 $\approx$1.1M
+boneS01 127 224 $\approx$3.3M del23 $\approx$8.4M $\approx$25.2M m14b 214 765 $\approx$1.7M
+Dubcova3 146 689 $\approx$1.7M rgg23 $\approx$8.4M $\approx$63.5M auto 448 695 $\approx$3.3M
+bmwcra_1 148 770 $\approx$5.2M del24 $\approx$16.8M $\approx$50.3M
+G2_circuit 150 102 288 286 rgg24 $\approx$16.8M $\approx$132.6M
+shipsec5 179 860 $\approx$5.0M europe_osm $\approx$50.9M $\approx$54.1M
+cont-300 180 895 448 799
 Benchmark Instances.
 To make our work more comparable, we choose the same benchmark instances as in [45, 16, 25, 44, 48].
-They can be seen in Table I.
+They can be seen in Table 1.
 The small instances come from the SuiteSparse Matrix Collection [12].
-These represent task graphs from real application.
+These represent task graphs from real applications.
 Six graphs were taken from Walshaws’s benchmark archive [46], which holds instances for benchmarking graph partitioning algorithms.
-The graphs af_shell9, thermal2, nlr are from the matrix and numerical section of the 10th DIMACS Implementation Challenge [5].
-The graphs deu and eur are large road networks of Germany and Europe [12].
+The graphs afshell9, thermal2, nlr are from the matrix and numerical section of the 10th DIMACS Implementation Challenge [6].
+The graphs deu and europe_osm are large road networks of Germany and Europe [13].
 Graphs rgg23 and rgg24 are random geometric graphs with $2^{23}$ and $2^{24}$ vertices respectively.
-Each vertex represents a random point in the unit square, and if their distance is below than $0.55\sqrt{\ln n/n}$ they are connected via an edge.
+Each vertex represents a random point in the unit square, and if their distance is less than $0.55\sqrt{\ln n/n}$ they are connected via an edge.
 Graphs del23 and del24 also have $2^{23}$ and $2^{24}$ vertices, with each vertex representing a point in the unit square.
-The edges are connected such that the graph is Delaunay triangulation.
-TABLE I: Benchmark instance properties.
-SuiteSparse Other
-Graph $|V|$ $|E|$ Graph $|V|$ $|E|$
-cop20k_A 99 843 1 262 244 af_shell9 $\approx$504K $\approx$8.5M
-2cubes_sphere 101 492 772 886 thermal2 $\approx$1.2M $\approx$3.7M
-thermomech_TC 102 158 304 700 nlr $\approx$4.2M $\approx$12.5M
-cfd2 123 440 1 482 229 deu $\approx$4.4M $\approx$5.5M
-boneS01 127 224 3 293 964 del23 $\approx$8.4M $\approx$25.2M
-Dubcova3 146 689 1 744 980 rgg23 $\approx$8.4M $\approx$63.5M
-bmwcra_1 148 770 5 247 616 del24 $\approx$16.7M $\approx$50.3M
-G2_circuit 150 102 288 286 rgg24 $\approx$16.7M $\approx$132.6M
-shipsec5 179 860 4 966 618 eur $\approx$18.0M $\approx$22.2M
-cont-300 180 895 448 799
-Walshaw
-598a 110 971 741 934 wave 156 317 1 059 331
-fe_ocean 143 437 409 593 m14b 214 765 1 679 018
-144 144 649 1 074 393 auto 448 695 3 314 611
-(a) Solution quality
-(b) Speedup
-Algorithm Phase Small Large GPU-HM I/O 24.9% 47.9% Jet 75.1% 52.1% GPU-HM-ultra I/O 4.8% 13.7% Jet 95.2% 86.3% GPU-IM I/O 29.3% 74.9% Heavy-Edge Matching 0.4% 1.1% Edge-Contraction 0.4% 4.7% KaFFPa 64.8% 6.2% Label Prop.
-+ Bal.
-5.0% 12.9% Uncoarsening 0.1% 0.2%
-(c) Runtime breakdown
-Figure 1: Comparison of solution quality (a) and speedup (b) of our proposed algorithms. Table (c) shows the runtime breakdown per phase.
+The edges are connected such that the graph is a Delaunay triangulation.
 Experimental Setup.
-As the hierarchy we choose $H=4:8:\{1,\ldots,6\}$ and as the distance we choose $D=1:10:100$.
+As the hierarchy, we choose $H=4:8:\{1,\ldots,6\}$ and as the distance, we choose $D=1:10:100$.
 The imbalance parameter is set to 3% ($\epsilon=0.03$).
-Each algorithm is run five times with different seeds and the running time and communication cost $J(C,D,\Pi)$ are averaged across the seeds.
+Each algorithm is run five times with different seeds and the running time and communication cost $J(\mathcal{C},\mathcal{D},\Pi)$ are averaged across the seeds.
+We do not include the time to initially read the graph from disk and constructing it on the CPU, and also not the time to write the partition to disk.
+These steps are conceptually the same for all algorithms.
 Performance Profiles.
-Performance Profiles [13] enable us to analyze the quality of multiple algorithms.
+Performance Profiles [14] enable us to analyze the quality of multiple algorithms across multiple instances.
 Let $\mathcal{A}$ be the set of all algorithms and let $\mathcal{I}$ be the set of all instances.
 The quality of algorithm $A\in\mathcal{A}$ on instance $I\in\mathcal{I}$ is $q_{A}(I)$ and $\text{Best}(I)=\min_{A\in\mathcal{A}}q_{A}(I)$ denotes the best quality across all algorithms on instance $I$.
-For each algorithm $A$, a performance plot shows the fraction of instances, the $y$-axis, for which $q_{A}(I)\leq\tau\text{Best}(I)$, where $\tau\geq 1$ is on the $x$-axis.
-For $\tau=1$ the plot shows the fraction of instances that each algorithm solved with the best solution.
+For each algorithm $A$, a performance profile shows the fraction of instances on the $y$-axis, for which $q_{A}(I)\leq\tau\text{Best}(I)$, where $\tau\geq 1$ is on the $x$-axis.
+For $\tau=1$, the profile shows the fraction of instances that each algorithm solved with the best solution.
 Algorithms that perform well have a high fraction of instances solved with minimal $\tau$ values, they are towards the top-left corner.
 Sub-optimal algorithms will have small fractions at minimal $\tau$ values, and only achieve higher fractions with greater $\tau$ values.
-V-A Results
-We will first analyze the results of our GPU-based hierarchical multisection (GPU-HM) and integrated mapping (GPU-IM) approach.
-Note that Jet [18], which we utilize in GPU-HM, has an ultra configuration which executes 18 refinement iterations instead of only one, to improve solution quality.
-We abbreviated our hierarchical multisection approach that uses Jet-ultra as GPU-HM-ultra.
-In the end, we will compare our methods against the established algorithms SharedMap [44] and IntegratedMapping [15], as they are CPU-based equivalent algorithms to hierarchical multisection and integrated mapping.
-Comparison.
-Figure 1 shows the performance profile (a) and the speedup (b) of our algorithms across all instances.
+5.1 Results
+First, we analyze our integrated mapping (GPU-IM) and hierarchical multisection (GPU-HM) algorithms.
+Jet [19], which is utilized in GPU-HM, also has an ultra configuration, which executes the complete refinement algorithm 18 times.
+This yields better solution quality at the cost of runtime.
+We abbreviate our approach that uses the ultra variant as GPU-HM-ultra.
+In the end, we compare our methods against the established algorithms SharedMap [45] and IntMap [16], which are CPU-based counterparts to hierarchical multisection and integrated mapping.
+Figure 1: Performance profile comparing the solution quality (left) and a speedup plot (right) of our proposed algorithms.
+5.2 Own Comparison
+Figure 1 shows the performance profile and the speedup of our algorithms across all instances.
 GPU-HM-ultra serves as a baseline to measure the speedup as it achieves the lowest communication cost.
-The default GPU-HM approach reaches a maximum speedup of $6.4$ with a geometric mean speedup of $4.4$.
+The default GPU-HM approach reaches a maximum speedup of $9.1\times$ with a geometric mean speedup of $6.5\times$.
 However, the speedup comes with an increase in communication cost.
-While the ultra variant finds the best solution on 94.794.7% of instances, GPU-HM does so on only 5.35.3%.
-On average, the communication cost of GPU-HM-ultra is just 0.10.1% above the best solution, whereas GPU-HM is 9.39.3% worse.
-GPU-IM is the fastest of our algorithms with a maximum speedup of $16$ and a geometric mean speedup of $5.6$.
-However, it is also the worst in terms of communication cost, as it did not achieve the lowest communication cost on any instance and has on average 17.417.4% higher communication cost than the best solution.
-Table 1c reports the percentage of runtime each algorithm spends in its major phases.
-Results are grouped by graph size: small graphs with fewer than one million vertices and large graphs with more than one million vertices.
-For the hierarchical multisection approach, most runtime is spent in the partitioning, which gets executed by Jet [18].
-Only for large graphs does I/O become significant, accounting for 47.947.9% of the runtime in the case of GPU-HM.
-The I/O phase includes reading the graph from disk, constructing its data structures, transferring it to the GPU and writing the final partition to disk.
-Among these steps, loading and constructing the graph dominate the I/O runtime.
-Overall, reducing the runtime of Jet or employing a faster GPU-based partitioner appears to be the most promising direction for improving performance further.
-The runtime distribution for GPU-IM follows a similar pattern.
-The greatest runtime percentages are in the phases that do not utilize the GPU, namely I/O and utilizing KaFFPa [40] for the initial partitioning.
-For smaller graphs, one could coarsen the graph further to partition a smaller graph, which would reduce the required time.
-However, this often leads to imbalanced solutions and worse communication cost.
-As later results will show, GPU-IM already lags behind other CPU-based algorithms in terms of solution quality.
-We therefore utilize GPU-IM as a very fast but low-accuracy algorithm.
-For larger graphs, the integrated mapping approach spends most of its runtime in the I/O phase.
-Figure 2: Comparison of speedup (left) and performance profile of achieved communication costs (right) of our proposed algorithms GPU-HM-ultra and GPU-IM to CPU-based algorithms SharedMap and IntegratedMapping.
-Comparison to CPU-based Algorithms.
-We compare our algorithms to the CPU-based solvers SharedMap [44] and IntegratedMapping [15], since they performed well in the experimental evaluation of [44].
+While the ultra variant finds the best solution on $95.3\%$ of instances, GPU-HM does so on only $4.7\%$.
+On average, the communication cost of GPU-HM-ultra is just $0.2\%$ above the best solution, whereas GPU-HM is $5.1\%$ worse.
+GPU-IM is the fastest of our algorithms with a maximum speedup of $150.1\times$ and a geometric mean speedup of $64.9\times$.
+However, it did not achieve the lowest communication cost on any instance and has, on average, $17.4\%$ higher communication cost than the best solution.
+Table 2 reports the average percentage of runtime GPU-IM spends in its major phases.
+The hierarchical multisection approaches spend more than $95\%$ of runtime in the partitioning, so we do not analyze it any further.
+The results in the table are grouped by graph size: small graphs with fewer than one million vertices and large graphs with more than one million vertices.
+For smaller graphs, the most time consuming phase is the refinement phase where the algorithm almost spends two-thirds of the runtime.
+Both the coarsening and the initial partitioning phase take about $13\%$ of runtime, and the other phases have a negligible impact on the running time.
+For larger graphs the distribution is more balanced.
+Notably, refinement and rebalancing is again the most time consuming phase, with about $45\%$ of total runtime.
+The initial partitioning now also becomes negligible, however coarsening and contraction take about $11\%$ each.
+Interestingly all miscellaneous operations, which also include uploading and downloading the initial graph and the resulting partition from the CPU to the GPU and vice versa, is the second greatest runtime consumption.
+This phase is mostly hardware dependent and cannot be sped up by further algorithm improvements.
+For the largest and smallest graphs, namely cop20k_A and europe_osm, the absolute running time of each phase is shown in milliseconds for the hierarchy $4:8:6$.
+Table 2: Runtime distribution across GPU-IM algorithm phases. The last two columns hold the absolute running times of each phase in milliseconds.
+Phase Small Large cop20k_A (ms) europe_osm (ms)
+Coarsening 13.02% 11.59% 4.351 41.020
+Contraction 3.49% 11.16% 1.010 38.694
+Init. Part. 13.85% 4.23% 11.116 7.244
+Uncontr. 0.14% 0.24% 0.046 1.523
+Refine + Reb. 65.22% 45.53% 24.359 116.598
+Misc 4.28% 27.24% 1.193 115.469
+Total 100% 100% 42.076 320.550
+5.3 CPU-based Comparison
+We compare our algorithms to the CPU-based solvers SharedMap [45] and IntMap [16], since they performed well in the experimental evaluation of [45].
 SharedMap employs the hierarchical multisection approach, just like GPU-HM.
-For partitioning, it internally uses KaFFPa [40] in either the Fast or Strong configuration, which trade speed for solution quality.
-We abbreviate the variants by SharedMap-F and SharedMap-S.
-IntegratedMapping, in contrast, minimizes $J(C,D,\Pi)$ via the multilevel scheme, making it conceptually similar to GPU-IM.
-It also provides a Fast and Strong configuration, which we denote by IntMap-F and IntMap-S.
-For comparison, we focus on GPU-HM-ultra, which achieves the highest solution quality, and GPU-IM, which provides the lowest runtime.
+It internally uses KaFFPa [41] for its graph partitioning.
+IntMap, in contrast, minimizes $J(\mathcal{C},\mathcal{D},\Pi)$ via the multilevel scheme, making it conceptually similar to GPU-IM.
+Both algorithms have a Fast and Strong configuration which we denote by the suffix -F and -S, respectively.
+These configurations trade speed for solution quality.
+For comparison, we focus on GPU-HM-ultra, which achieves the highest solution quality, and GPU-IM, which provides the lowest running time.
 GPU-HM is omitted for better visual clarity, as its performance lies between the two.
-Figure 2 shows the speedup over SharedMap-S (left) and the performance profile (right).
-Among all algorithms, SharedMap-S achieves the highest solution quality, obtaining the best solution on 125 instances (83.383.3%).
+Figure 2 shows the speedup over SharedMap-S and the performance profile.
+Among all algorithms, SharedMap-S achieves the highest solution quality, obtaining the best solution on 124 instances ($82.7\%$).
 However, this comes at the cost of runtime, as SharedMap-S is also by far the slowest algorithm.
-GPU-HM-ultra finds the best solution on the remaining 25 instances (16.716.7%).
+GPU-HM-ultra finds the best solution on the remaining 26 instances ($17.3\%$).
 All other algorithms are not able to determine a best solution on any instance.
-Ranked by average additional cost over the best solution, the algorithms are ordered as follows: SharedMap-S (0.020.02%), GPU-HM-ultra (11.511.5%), IntMap-S (14.514.5%), IntMap-F (21.121.1%), GPU-IM (30.730.7%) and SharedMap-F (31.731.7%).
-GPU-IM and SharedMap-F are at most $2.42\times$ and $2.66\times$ worse than the best-known solution, respectively.
-In terms of runtime, GPU-IM is the fastest algorithm overall, achieving a maximum speedup of $598.5$ and a geometric mean speedup of $77.7$ compared to SharedMap-S.
-While GPU-HM-ultra achieves a maximum speedup of $377$, its geometric mean speedup is only $13.8$.
-On many instances, it is slower than IntMap-F/S and SharedMap-F, largely due to the partitioning phase being the bottleneck (see Table 1c).
-The remaining mean speedups are $28.4$ for SharedMap-F, $23.8$ for IntMap-F, and $10.4$ for IntMap-S.
-Our algorithms have two clear equal CPU-based counterparts.
-GPU-HM-ultra slightly outperforms IntMap-S in terms of solution quality.
-Its geometric mean speedup of $13.8$ also exceeds the $10.4$ of IntMap-S, although it is not faster across all instances.
-Specifically, IntMap-S is faster on 7171% of small instances (speedup of $1.4$), whereas GPU-HM-ultra dominates on 9696% of the large instances (speedup of $5.6$).
-Similarly, GPU-IM is comparable to SharedMap-F in terms of solution quality.
-On small graphs, it achieves on average 2.62.6% lower communication costs, while on larger graphs it performs about 2.32.3% worse.
-The geometric mean speedup of GPU-IM over SharedMap-F is $1.4$ and $10.6$ for small and large graphs respectively.
-In summary, GPU-HM-ultra outperforms the CPU-based algorithm IntMap-S in both solution quality and running time.
-However, it cannot reach the state-of-the-art solution quality of SharedMap-S.
-Nevertheless, GPU-HM-ultra is on average an order of magnitude faster.
-GPU-IM can be almost up to $600$ times faster than SharedMap-S, but its solution quality is on average 30.730.7% higher than the best-known solution.
-Comparison to GPU-based Algorithms.
+Ranked by average additional cost over the best solution, the algorithms are ordered as follows: SharedMap-S ($0.2\%$), GPU-HM-ultra ($12.2\%$), IntMap-S ($14.4\%$), IntMap-F ($20.9\%$), SharedMap-F ($30.8\%$), and GPU-IM ($33.1\%$).
+SharedMap-F and GPU-IM are at most $1.99\times$ and $3.07\times$ worse than the best-known solution, respectively.
+Figure 2: Comparison of the speedup (left) and the performance profile of achieved communication costs (right) of our proposed algorithms GPU-HM-ultra and GPU-IM to CPU-based algorithms SharedMap [45] and IntMap [16].
+In terms of runtime, GPU-IM is the best algorithm overall, achieving a maximum speedup of $12376.9\times$ and a geometric mean speedup of $1454.6\times$ compared to SharedMap-S.
+In fact, it is the fastest algorithm on all 150 instances.
+While GPU-HM-ultra achieves a maximum speedup of $934.7\times$, its geometric mean speedup is only $22.4\times$.
+The remaining mean speedups are $42.7\times$ for SharedMap-F, $36.7\times$ for IntMap-F, and $11.7\times$ for IntMap-S.
+Our algorithms have two clear CPU-based counterparts.
+GPU-HM-ultra outperforms IntMap-S in terms of solution quality and speed.
+It finds better solutions on 117 out of 150 instances ($78.0\%$) and is also faster on 93 instances ($62.0\%$).
+When split into small and large graphs, with fewer and more than one million vertices respectively, GPU-HM-ultra is only faster on 45 out of 102 small instances, but faster on all 48 large instances with a mean speedup of $12.6\times$.
+When inspecting the communication cost, GPU-HM-ultra has better solution quality on 95 small instances, while only being better on 22 large instances.
+Similarly, GPU-IM’s solution quality lies between SharedMap-F and IntMap-F.
+On smaller graphs GPU-IM finds better solutions than SharedMap-F on 70 instances, but only on 5 instances when compared to IntMap-F.
+For larger graphs this ratio worsens as GPU-IM only finds better solutions on 14 and 10 instances when compared to SharedMap-F and IntMap-F respectively.
+However, GPU-IM is much faster than both algorithms with a mean speedup of $55.8\times$ and a max speedup of $221.6\times$ over SharedMap-F.
+Compared to IntMap-F, the mean speedup is $70.8\times$ and the max speedup is $243.1\times$.
+In summary, GPU-HM-ultra outperforms IntMap-S in both solution quality and running time.
+While being about $20$ times faster than SharedMap-S, it cannot reach its state-of-the-art solution quality.
+GPU-IM achieves a mean speedup of $1454\times$ compared to SharedMap-S, but its communication cost is on average $33.1\%$ higher.
+Nonetheless, its quality is comparable to IntMap-F and SharedMap-F while being about $60$ times faster.
+5.4 GPU-based Comparison
 While no other GPU-based algorithms currently exist for solving the GPMP, we can compare our algorithms to Jet [19].
-Note that Jet solves the graph partitioning problem, which corresponds to using a distance vector of $D=1:\ldots:1$.
+Note that Jet solves the graph partitioning problem, which corresponds to using the distance vector $D=1:\ldots:1$.
 Since we test the distance $D=1:10:100$, Jet’s resulting partitions are unfit and the corresponding communication costs are worse than those of our algorithms.
 In terms of runtime, however, Jet is comparable to GPU-IM as they both follow the multilevel scheme.
-We base our comparison on the large graph instances ($\geq 1$ million vertices).
-Using GPU-IM as the baseline for runtime and communication costs, we observe that Jet and Jet-ultra incur additional communication costs of 42.142.1% and 38.838.8% on average respectively.
-In terms of runtime, Jet-ultra is comparable to GPU-IM, with a geometric mean speedup of $1.03$, whereas Jet achieves an average speedup of $1.33$.
-We attribute this performance gap primarily to the 4040% matching requirement of our algorithm, compared to the 7575% in Jet.
-This allows Jet to produce smaller coarse graphs in fewer levels, and thereby reduce runtime.
-While we could tune our constants for more speed, we observed that it negatively impacted solution quality.
-Our algorithms GPU-HM and GPU-HM-ultra further improve solution quality but are worse in runtime than GPU-IM (see Figure 1), which translates to this GPU comparison.
-VI Conclusion
+Jet’s determined partitions incur on average $45.3\%$ additional communication cost over those by GPU-IM.
+Compared to SharedMap-S this is an average increase of $90.3\%$.
+As explained it is expected that the produced partitions are unfit for process mapping, which highlights the need for a dedicated solver.
+Note that the quality of the ultra configuration is even worse.
+This can be explained because decreasing the edge-cut does not necessarily decrease the communication cost.
+Trading two cut edges for one cut edge is always beneficial (assuming uniform edge weights) when optimizing the edge-cut, however this is not inherently true for process mapping.
+Trading two cut edges on neighboring PEs for one cut edge that connects two islands has an overall detrimental effect.
+In terms of runtime, GPU-IM is slightly faster than Jet.
+On small graphs, with fewer than one million vertices, the geometric mean speedup is $1.43\times$, with a minimum and maximum speedup of $0.21\times$ and $1.95\times$.
+For larger graphs, with more than one million vertices, the mean speedup increases to $1.56\times$.
+The minimum and maximum speedup are $1.21\times$ and $2.22\times$.
+We primarily attribute the small speed gains to the use of the extended CSR format and miscellaneous technical improvements.
+In summary, GPU-IM has a geometric mean speedup of $1.47\times$ over Jet and therefore delivers equal speed to current state-of-the-art GPU-based graph partitioners.
+6 Conclusion
 Process mapping assigns vertices of a task graph to the processing elements (PEs) of a supercomputer to balance workload and minimize communication costs.
 Motivated by the success of GPU-based graph partitioners, we proposed two GPU-accelerated algorithms for this problem.
-The first, GPU-HM, applies hierarchical multisection [47] using the GPU-based partitioner Jet [18].
+The first, GPU-HM, applies hierarchical multisection [48] using a reimplementation of the GPU-based graph partitioner Jet [19].
 The second, GPU-IM, integrates process mapping directly into a multilevel framework, adapting Jet’s label propagation and rebalancing.
-GPU-HM-ultra achieves up to $300\times$ speedups and is on average a magnitude faster than the state-of-the-art SharedMap [44], with only 11.511.5% higher communication costs and therefore outperforming IntMap-S [15].
-GPU-IM delivers extreme speedups (mean $77.7\times$, max $598.5\times$) at lower solution quality, however it still outperforms the Fast configuration of SharedMap.
-Future work will focus on improving solution quality without sacrificing performance e.g., enhancing label propagation filters, refining coarsening and initial partitioning, and exploring more efficient GPU-based partitioners.
+GPU-HM-ultra achieves up to $934\times$ speedups and is on average $20$ times faster than the state-of-the-art algorithm SharedMap [45], with only $12.0\%$ higher communication costs and thereby outperforming IntMap-S [16].
+GPU-IM delivers extreme speedups (mean $1454\times$, max $12376\times$) at lower solution quality, however it is still comparable the Fast configuration of both SharedMap [45] and IntMap [16].
+Future work will focus on improving solution quality without sacrificing performance, e.g., enhancing label propagation filters and refining initial partitioning.